### v2

# Enhancing Clinical Trial Patient Matching through Knowledge Augmentation and Reasoning with Multi-Agents

Abstract.
Matching patients effectively and efficiently for clinical trials is a significant challenge due to the complexity and variability of patient profiles and trial criteria. This paper introduces Multi-Agents for Knowledge Augmentation and Reasoning (MAKAR), a novel framework that enhances patient-trial matching by integrating domain-specific knowledge with structured reasoning. MAKAR consists of two key modules: the Augmentation Module and the Reasoning Module. The Augmentation Module enriches trial criteria by incorporating detailed explanations of relevant concepts to ensure clarity and completeness. The Reasoning Module then evaluates each health condition, following a structured, step-wise approach to determine eligibility and make the final matching decision. This paper enhances patient-trial matching by leveraging the agency and reasoning capabilities of large language models (LLMs) through automated agent interactions, including collaboration, critique, and navigation. Experimental results on a public dataset demonstrate that our framework surpasses existing benchmarks, achieving up to an 8% improvement in accuracy for specific criteria. Furthermore, in a real-world offline test, MAKAR achieved a 100% accuracy. These findings show MAKAR’s potential as a scalable and robust solution for clinical trial patient matching.
Patient Matching, Multi-agent Systems, Knowledge Augmentation, Reasoning LLMs
ccs: Computing methodologies Multi-agent systemsccs: Applied computing Health care information systems
1. Introduction
Patient matching for clinical trials refers to the process of identifying and enrolling participants whose health profiles meet the specific eligibility criteria of a given clinical study. Efficient patient matching is crucial to accelerating the drug development process and reducing the overall cost of clinical trials. It also provides an opportunity for participants to access experimental treatments that could potentially improve their health outcomes and, in some cases, transform their quality of life.
Patient matching remains a challenging task despite its critical role in clinical trials. Traditional automated methods, which are mostly rule-based, depend heavily on hand-crafted rules and are often time-consuming (Brøgger-Mikkelsen et al., 2020). These approaches are also highly sensitive to variations in how patient and trial data are expressed in natural language across different sources. Simple machine learning models often struggle with the load of information and the low density of relevant sentences in lengthy electronic health record(EHR) data. They usually underperform when dealing with complex eligibility criteria or inconsistent data formats, resulting in inefficiencies and mismatches that can negatively impact trial outcomes (Kadam et al., 2016).
The emergence of of LLMs offers a promising opportunity to improve patient matching. LLMs, such as GPT-4 (Achiam et al., 2023), usually possess a large text window and advanced natural language understanding capabilities (Karanikolas et al., 2023), which enable them process complex trial criteria and patient data effectively. However, their general knowledge may not always be sufficient for knowledge-intensive tasks like patient matching in clinical trials, where domain-specific expertise is required. The knowledge gap inherent in LLMs and misssing information in the pre-defined criteria can limit their performance, highlighting the need for LLM augmentation of specialized medical knowledge (Shuster et al., 2021). While research indicates that LLMs do encode medical knowledge in a parameterized form (Singhal et al., 2023), this information can be outdated or misaligned with the special needs of certain clinical trials.
Figure 1. The workflow of MAKAR framework
To address these limitations, we present a new framework called: Multi-Agents for Knowledge Augmentation and Reasoning (MAKAR). MAKAR is a novel framework designed to enhance the patient matching process. The framework consists of two core modules: the Augmentation Module and the Reasoning Module. The Augmentation Module enhances trial criteria by providing detailed explanations and includes three key components: a Router Agent, multiple Augmentation Agents, and a Critic Agent. The Reasoning Module systematically evaluates each criterion-related condition in a step-by-step manner to determine eligibility and make the final matching decision. It comprises a Reasoning Agent and a Supervision Matching Agent. By leveraging this multi-agent framework, MAKAR improves the performance on patient-trial matching by dynamically incorporating domain-specific knowledge and step-wise reasoning. Our experimental results demonstrate the effectiveness of MAKAR. On a public dataset, our framework achieves an average improvement of 3% over benchmark methods, with up to an 8% improvement for complex criteria. Furthermore, in an offline real-world data test, MAKAR outperformed the state-of-the-art closed-source reasoning LLMs and achieved 100% accuracy, showing its reliability in practical applications. Our key contributions are as follows:
  • We propose an automated method to refine and enhance trial criteria, making them more interpretable and suitable for AI-driven patient-trial matching.
  • We introduce a multi-agent framework that combines augmentation and reasoning, achieving strong performance across both public and real-world datasets.
  • We identify a gap in existing approaches and highlight the importance of explicitly incorporating well-structured, detailed explanations of related concepts to improve LLM-driven clinical reasoning.
2. Related Work
2.1. Clinical Trial Patient Matching
Traditional patient matching systems rely on query-based structured databases, such as Criteria2Query (Yuan et al., 2019), or handcrafted rule-based approaches like RBC (Oleynik et al., 2019). However, these methods struggle to handle unstructured free-text data and are highly sensitive to variations in formatting and verbal expressions. Unfortunately, a significant portion of EHRs consists of free-text, including clinical notes, surgical records, radiology reports, pathology reports and so on (Sun et al., 2018). Thus, the effectiveness and scalability of traditional methods are significantly limited, making them unsuitable for future patient matching tasks.
Recent advancements in LLMs have shown remarkable capabilities in directly processing free-text clinical trial criteria, offering scalable solutions without the need for extensive rule-based preprocessing. Many Studies have applied these models to patient-centric trial matching tasks using real-world data. For instance, Nievas et al. (2024) fine-tuned the open-source Llama 2 (Touvron et al., 2023) to automate patient matching. However, fine-tuning LLMs is computationally expensive, prompting alternative approaches that leverage frozen models. Jin et al. (2023) introduced TrialGPT, a patient-centric matching framework that assesses patient eligibility on a criterion-by-criterion basis, filtering out ineligible trials and ranking potentially eligible ones. Their results indicated a strong correlation between TrialGPT’s predictions and expert clinician judgments, achieving accuracy at the criterion level similar to human experts. Similarly, Gupta et al. (2024) proposed OncoLLM, an end-to-end pipeline that converts trial criteria into conversational prompts and uses a Chain-of-Thought (CoT) (Wei et al., 2022) reasoning strategy. OncoLLM outperforms GPT-3.5 while underperforms GPT-4, despite its smaller model size.
In contrast, other studies have focused on trial-centric patient matching, typically using public deidentified datasets due to HIPAA-related data constraints. Wornow et al. (2024) developed a LLM-based system with a two-stage retrieval pipeline to identify suitable patients for trials using a zero-shot prompt approach. It reduces token processing costs while outperforming existing state-of-the-art models on the n2c2 2018 cohort selection benchmark (Stubbs et al., 2019). Likewise, Beattie et al. (2024) applied a CoT strategy in a Retrieval-Augmented Generation (RAG) (Lewis et al., 2020) system using GPT-3.5 Turbo and GPT-4, demonstrating results comparable to those of Wornow et al. (2024).
These works collectively highlight the potential of LLMs in clinical trial matching. However, existing approaches directly employ LLMs without verifying whether the trial criteria are sufficiently understood by the model (Webson and Pavlick, 2021) or whether there are gaps in the model’s knowledge (Lewis et al., 2020). Our work fits within the trial-centric patient matching literature and makes a novel contribution by automatically identifying and addressing these gaps, thereby enhancing the accuracy patient matching.
2.2. LLM-based Multi-Agents (LLM-MA)
LLM-MA have seen rapid development across various domains, driven by the promising potential of LLMs as autonomous agents.  Guo et al. (2024) provide a comprehensive overview of recent advancements in LLM-MA applications, categorizing them into two primary areas: Problem-Solving and World Simulation. Problem-solving LLM-MA systems explore the capabilities of agents with specialized expertise, enabling them to outperform single-agent LLM systems through collaborative efforts. For instance,  Wang et al. (2024) proposed Mixture-of-Agents, a layered multi-agent architecture that achieves state-of-the-art performance in benchmarks like AlpacaEval 2.0 (Dubois et al., 2024).
In the healthcare domain, the application of LLM-MA systems is rapidly gaining traction. Yue et al.(Yue and Fu, 2024) introduce ClinicalAgent, a multi-agent architecture powered by LLM reasoning for multi-agent collaboration, achieving higher precision in predicting clinical trial outcomes. Similarly, Liao et al.(Liao et al., 2024) propose REFLECTool, a multi-agent framework that enables agents to learn how to select domain-specific tools, thereby extend their capabilities. TriageAgent(Lu et al., 2024) leverages LLMs for role-playing interactions, integrating self-confidence assessments and early-stopping mechanisms in multi-round discussions to improve document reasoning and classification accuracy for triage tasks. Tang et al.(Tang et al., 2023) use multiple LLM-based agents as domain experts to collaboratively analyze medical reports. And Li et al. (Li et al., 2024) further introduce the multi-agent collaboration framework to handle multimodal tasks.
To the best of our knowledge, this study is the first to apply an LLM-MA framework for cohort selection in clinical trials. Our approach explore the agency and reasoning capabilities of LLM-MA to dynamically enhance patient-trial matching with tailored knowledge. Our approach allows a better contextual understanding of patient profiles and lead to improved matching performance and more informed decision-making in clinical trial recruitment.
Criteria Accuracy Precision Recall F1-score
Zero-shot CoT MAKAR Zero-shot CoT MAKAR Zero-shot CoT MAKAR Zero-shot CoT MAKAR
ABDOMINAL 0.858 0.868 0.847 0.893 0.897 0.944 0.701 0.729 0.626 0.785 0.804 0.753
ADVANCED-CAD 0.854 0.854 0.868 0.895 0.872 0.929 0.853 0.882 0.841 0.873 0.877 0.883
ALCOHOL-ABUSE 0.969 0.969 0.976 0.529 0.545 0.667 0.900 0.600 0.600 0.667 0.571 0.632
ASP-FOR-MI 0.861 0.875 0.885 0.899 0.908 0.919 0.930 0.939 0.939 0.915 0.923 0.929
CREATININE 0.802 0.802 0.885 0.664 0.667 0.812 0.934 0.925 0.896 0.776 0.775 0.852
DIETSUPP-2MOS 0.764 0.781 0.785 0.704 0.717 0.721 0.940 0.953 0.953 0.805 0.818 0.821
DRUG-ABUSE 0.941 0.920 0.993 0.469 0.395 0.882 1.000 1.000 1.000 0.638 0.566 0.938
ENGLISH 0.976 0.986 0.983 0.985 0.989 0.996 0.989 0.996 0.985 0.987 0.992 0.991
HBA1C 0.840 0.830 0.878 0.722 0.702 0.807 0.892 0.902 0.863 0.798 0.790 0.834
KETO-1YR 0.993 0.990 0.993 - - - - - - - - -
MAJOR-DIABETES 0.833 0.826 0.847 0.851 0.831 0.873 0.840 0.853 0.840 0.845 0.842 0.856
MAKES-DECISIONS 0.920 0.924 0.931 0.985 0.985 0.996 0.931 0.935 0.931 0.957 0.959 0.963
MI-6MOS 0.885 0.872 0.927 0.434 0.396 0.571 0.885 0.808 0.769 0.582 0.532 0.656
Average 0.884 0.884 0.908 0.750 0.740 0.841 0.899 0.876 0.853 0.801 0.786 0.841
Table 1. Criterion-Level Comparison Across Different Methods for N2C2
3. Methods
The proposed framework, shown in Figure 1, consists of two core modules: the Augmentation Module and the Reasoning Module. The augmentation module is designed to generate criteria with comprehensive explanations of all relevant concepts, while the reasoning module evaluates each condition within the criteria to determine eligibility and make the final matching decision. Additionally, the framework contains five interconnected components, each outlined below.
Router Agent In the MAKAR framework, multiple specialized augmentation agents are available to augment the criterion with domain-specific knowledge. The Router Agent determines the most suitable augmentation agent for the task and routes the workflow accordingly. Inspired by SELFREF (Kadavath et al., 2022) and MOREINFO (Feng et al., 2023), we adopt a prompt-based self-probing strategy. If the agent determines that the original criterion is detailed and understandable enough for other agents, the criterion will be directly forwarded to the Reasoning Module. Otherwise, the criterion is forwarded to the Knowledge Augmentation Agent chosen by the Router Agent for further processing.
Knowledge Augmentation Agent: This component augments the criterion with relevant, domain-specific information from different source. For instance, the Retrieval Agent (Lewis et al., 2020) retrieves information from indexed medical databases, the Online Search Agent gathers supplementary knowledge from online sources, and the Self-Augment Agent leverages the LLM’s intrinsic knowledge to refine the original criterion.
Critic Agent: Acting as a supervisor, this agent ensures that the augmented criterion is aligned with the original instructions. For instance, the original criterion must be preserved verbatim. Additionally, the critic agent verifies that the augmented output is well-structured (e.g., concepts are clearly itemized) to facilitate further step-wise reasoning. This critic mechanism are designed to improve the framework’s ability to follow instructions accurately while maintaining consistency and clarity.
Reasoning Agent With the augmented criterion and patient EHR data, the Reasoning Agent evaluates patient eligibility by assessing for each concepts mentioned in the augmented criterion. It generates a detailed reasoning summary, including assessments for all relevant health conditions, and forwards it to the Supervision Matching Agent.
Matching Agent: The Matching Agent determines patient eligibility by reviewing the reasoning assessment alongside the patient’s EHR data, making a final judgment in a zero-shot manner.
4. Evaluation & Results
Due to the sensitive nature of patient data, which is strictly protected under The Health Insurance Portability and Accountability Act (HIPAA), and the compliance requirements for processing detail clinical trial documents, full-scale online deployment is not currently feasible—particularly due to concerns about data security when using closed-source LLMs. Instead, to evaluate the effectiveness of our methods, we conducted two experiments: (1) an assessment using a public dataset and (2) a limited-scale offline evaluation with deidentified real-world data.
4.1. Data
The public dataset we use is from Track 1 of the 2018 n2c2 challenge (Stubbs et al., 2019), which is commonly used for evaluating patient matching in clinical trials. The dataset consists of 288 de-identified patients, each having between 2 to 5 unstructured longitudinal clinical notes in American English. The entire corpus contains 781,006 tokens, with an average of 2,711 tokens per patient. The challenge simulates a synthetic clinical trial with 13 predefined inclusion criteria, resulting in 3744 patient-criteria pairs in total. The detailed criteria can be found in Table 6. For each patient-criteria pair, a binary label is provided, indicating whether the patient ”MET” or ”NOT MET” the criterion. Two annotators with medical background labeled these data. The disagreements between the two annotators were resolved by an independent MD to establish the final ground truth labels.
The criteria in the n2c2 dataset are relatively straightforward and the patient records are well-structured, making them easy to follow. To evaluate the potential of our proposed method in a more complex scenario, we constructed a real-world dataset. Specifically, we sampled 30 eligibility criteria—15 inclusion and 15 exclusion—from ongoing breast cancer trials listed on ClinicalTrials.gov. We then manually selected 10 patients who have been diagnosed with breast cancer from our in-house EHR system. For each patient, we extracted their most recent deidentified clinical visit record. An expert subsequently asseessed whether each patient met each criterion, resulting in 300 labeled criterion-patient pairs. Notably, if a patient’s information was insufficient to determine eligibility for a given criterion, they were labeled as ”NOT MET” to ensure a conservative patient selection. We listed the sampled criterion in Table  9
4.2. Baselines and Metrics
Previous research has demonstrated the superiority of LLM-based methods over traditional rule-based systems; hence, this study focuses exclusively on LLM-based benchmarks. We include two approaches in Experiment 1: the zero-shot method with hand-crafted criteria (Wornow et al., 2024), as shown in Table LABEL:table:redefined_criteria, and a Chain-of-Thought (CoT) approach (Beattie et al., 2024) using the original criteria. In addtion, we also include o3-mini from OpenAI as the benchmark in Experiment 2 since reasoning LLMs have shown significant improvement on complex clinical scenarios(Xie et al., 2024). In line with prior literature, we report accuracy, precision, recall, and F1-score for this binary classification task.
Previous studies using n2c2 dataset have primarily focused on evaluating classification performance at the individual criterion level, often overlooking the overall patient-trial assessment. In this study, we also examine patient-trial selection performance in Experiment 1. Specifically, a patient is eligible for the trial only if all inclusion criteria are met. However, no patient in the n2c2 dataset satisfied all 13 criteria. To mitigate this, we synthesize a trial with selected criteria that had at least 100 patients meeting them, resulting in 28 eligible patients from a total cohort of 288. We then evaluated our method’s on this new synthetic trial in identifying eligible patients and reported the findings.
4.3. Experiment Settings
Our experiments were conducted on an Amazon EC2 g5.16xlarge instance, equipped with a A10G GPU with 24 GB of memory. Unless otherwise specified, all agents in the framework are powered by GPT-4o.
For external knowledge augmentation, the Online Search Agent uses the Perplexity AI search engine to fetch up-to-date information from the web. The Retrieval Agent uses all-MiniLM-L12-v2, a pre-trained checkpoint from Sentence-Transformer (Reimers, 2019), as its encoder to construct the vector database of MedDRA and FDALabel.
Metric Zero-shot CoT MAKAR
Accuracy 0.8924 0.9201 0.9306
Precision 0.4615 0.5676 0.6333
Recall 0.6429 0.7500 0.6786
F1-Score 0.5373 0.6462 0.6552
Table 2. Trial-Level Comparison Across Different Methods
4.4. Results of Experiment 1
4.4.1. Criterion-Level Performance:
Table 1 shows the performance of three approaches—Zero-shot,CoT, MAKAR—across multiple individual criteria. Our results indicate that while the CoT method slightly outperforms Zero-shot in Accuracy and Recall, MAKAR consistently achieves higher scores, particularly in Accuracy (0.909) and Precision(0.822) on average. The F1-score of MAKAR (0.828) suggests a balanced performance across criteria, although it falls slightly behind in a few specific cases such as ”ABDOMINAL” and ”English”.
The MAKAR-augmented criteria, presented in Table LABEL:tab:aug_criteria, demonstrate significant improvements over the baseline human-redefined criteria outlined in Table LABEL:table:redefined_criteria from Wornow et al. (2024). These enhancements primarily focus on providing explicit clarifications and detailed expansions of key concepts, enabling LLMs to better align with trial requirements. For example, MAKAR supplements normal ranges for criteria like ”HBA1C” and ”CREATININE” across diverse demographic groups. And they are also the criteria that benefit the most from MAKAR’s augmentation. Additionally, MAKAR provides a more comprehensive definition for criteria like ”MAJOR-DIABETES,” and use more widely recognized terminology, ”heart attack”, to explain clinical terms ”Myocardial Infarction (MI)”. These refinements improve the ability of LLMs to match patients from clinical notes written in a variety of styles, resulting in notable performance gains for these criteria.
4.4.2. Trial-Level Performance
The result of trial-levle analysis can be found in Table 2. At the trial level, where a patient’s eligibility is determined by satisfying all selected criteria, MAKAR outperforms both Zero-shot and CoT. MAKAR achieves the highest Accuracy (0.9306), Precision (0.6333)and F1-score (0.6552) but underperform CoT on Recall. The trial-level results further validate MAKAR’s capability in maintaining a higher balance between identifying eligible patients and minimizing patient-filtering load
Inclusion Criteria
Metric GPT-4o o1 o3-mini CoT MAKAR
Accuracy 0.9000 0.9333 0.8533 0.9667 1.000
Precision 0.7884 0.8407 0.7211 0.9385 1.000
Recall 0.9020 0.9413 0.8151 0.9208 1.000
F1-Score 0.8281 0.8800 0.7516 0.9294 1.000
Exclusion Criteria
Metric GPT-4o o1 o3-mini CoT MAKARCorr
Accuracy 0.9000 0.9267 0.9133 0.9467 1.000
Precision 0.8693 0.8987 0.8840 0.9265 1.000
Recall 0.9090 0.9416 0.9253 0.9487 1.000
F1-Score 0.8845 0.9153 0.8999 0.9365 1.000
Table 3. Criterion-Level Comparison Across Different Methods in Experment 2
4.5. Results of Experiment 2
4.5.1. Results
Table 3 presents the evaluation results of our method on the in-house real-world dataset. We compared its performance against zero-shot prompting with GPT-4o, zero-shot prompting with two reasoning models (OpenAI-o3-mini and OpenAI-o1) and CoT strategy with GPT-4o. Our method achieves perfect scores across all four metrics.
Notably, for exclusion criteria, we report results for a corrected version of MAKAR (MAKARCorr). In the initial version, the model was not explicitly instructed on whether a given criterion was for inclusion or exclusion. Surprisingly, it was able to infer some exclusion criteria but focused on determining a patient’s overall eligibility for a trial rather than assessing whether a specific criterion was met. For instance, if a patient was 75 years old, they should have been marked as ”MET” for an exclusion criterion of ”greater than 70 years old”. While MAKAR correctly identified the age range and provided a logical reasoning process, it mistakenly output ”Not MET”—interpreting the task as determining overall trial eligibility rather than verifying the criterion itself. The corrected version addresses this by explicitly instructing the model to evaluate each criterion independently rather than making a broader eligibility judgment.
MAKAR outperforms all the benchmarks and the CoT follows closely, achieving 96.67% accuracy and 92.94% F1-score for inclusion criteria and 94.67% accuracy with 93.65% F1-score for exclusion criteria. o3-mini, a smaller model, is outperformed by o1 but shows competitive performance on exclusion criteria with zero-shot method with GPT-4o. Interestingly, Experiment 2 yields better overall performance than Experiment 1, despite its more knowledge-intensive criteria. We hypothesize that this is due to differences in the labeling process—the n2c2 challenge involves more subjective judgment, while our setup explicitly defines insufficient information as ”NOT MET”.
4.5.2. Error Analysis
Despite MAKAR achieving 100% accuracy in the second experiment, we manually checked the reasoning results of all 300 patient-criterion pair and identified two reasoning errors that did not affect the final output.
First, in Exclusion Criterion 13 (Patients with ongoing indications for cardioprotective medication—ACE inhibitors, ARBs, and/or beta-blockers), MAKAR reasoned: ”The patient is on metoprolol, a beta-blocker, suggesting some ongoing indication for its use. However, without explicit information on the specific condition necessitating this medication (e.g., arrhythmia, prophylactic use post-cardiac event other than MI, or another reason), the justification remains uncertain.” This reasoning was overly conservative, but the Supervision Matching Agent correctly classified it as MET, aligning with expert judgment.
Second, in Inclusion Criterion 5 (Histologically confirmed breast cancer with an invasive component $\geq$ 20 mm and/or morphologically confirmed spread to regional lymph nodes(stage cT2-cT4 with any cN, or cN1-cN3 with any cT).), MAKAR reasoned: ”The documentation mentions a suspicious breast nodule at the right breast (1.4 × 1.5 × 0.5 cm), which measures more than 20 mm in one dimension.” This was an obvious error, as the maximum dimension stated is only 1.5 cm. The mistake may come from over-reasoning, possibly computing the largest possible diagonal as $\sqrt{1.4^{2}+1.5^{2}+0.5^{2}}=2.1cm$. However, the final decision remained correct since the criterion lymph node involvement still holds.
4.6. Ablation
In this section, we validate the contribution of each module in our framework. First, we remove the knowledge augmentation module to assess the reasoning module (MAKA-reason) using the original criteria, which essentially mirrors the CoT approach in Table 3. Next, we remove the reasoning module and evaluate the knowledge augmentation module (MAKA-aug) independently by performing zero-shot matching with the augmented criteria. The results, presented in Table 4, indicate that removing the reasoning module has a greater negative impact on inclusion criteria matching (F1-score: 88.37% vs. 92.94%), while removing the augmentation module has a larger effect on exclusion criteria (F1-score: 93.65% vs. 94.41%).
Although we did not involve any reasoning-LLM-based agents, our approach still shows impressive reasoning capabilities. This can be attributed to the well-structured augmented criteria and the CoT strategy employed by the reasoning agent. For instance, when evaluating whether a patient meets the Inclusion Criterion 14 (Postmenopausal hormone receptor-positive patients), MAKA reasoned: ”…the record does not explicitly state whether the breast cancer was hormone receptor positive. However, the use of tamoxifen, which is specifically prescribed for estrogen receptor-positive breast cancer by blocking estrogen receptors, strongly implies that her breast cancer was hormone receptor positive.” Both MAKA and MAKA-reason successfully made the correct match, while MAKA-aug failed to capture this implicit relationship.
Additionally, we also observed cases where knowledge augmentation improved performance. For example, in Exclusion Criterion 7, it mentions: ”…all histological lesions were HER2 1+ or 2+, not detected by FISH amplification, and HR positive according to ASCO guidelines”. MAKA-aug expanded the definitions from American Society of Clinical Oncology(ASCO) guidelines, made correct matches for all 10 patients. In contrast, MAKA-reason correctly identified only 8 out of 10.
Inclusion Criteria
Metric MAKAR-reason MAKAR-aug MAKAR
Accuracy 0.9667 0.9333 1.000
Precision 0.9385 0.8387 1.000
Recall 0.9208 0.9612 1.000
F1-Score 0.9294 0.8837 1.000
Exclusion Criteria
Metric MAKAR-reason MAKAR-aug MAKARCorr
Accuracy 0.9467 0.9533 1.000
Precision 0.9265 0.9360 1.000
Recall 0.9487 0.9534 1.000
F1-Score 0.9365 0.9441 1.000
Table 4. Results of Ablation Analysis
5. Discussion
One interesting observation from our experiments is that the Router Agent consistently determined that all criteria could be augmented by the Self-Augmentation Agent, without requiring external retrieval or online search augmentation. This was true even for breast cancer-related criteria, which are typically complex and challenging for individuals without professional training.
This relates to a limitation of our study, which comes from the dataset used for evaluation. Both the n2c2 dataset and the real-world dataset are relatively small and contain criteria that are not particularly challenging for LLM-based processing, allowing all benchmark methods to achieve strong performance. While MAKAR is designed as a general framework capable of handling both simple and complex criteria, these dataset constraints limit our ability to explore the knowledge boundaries of LLMs and fully assess the framework’s generalizability across a wider range of clinical trial requirements.
Moreover, this observation raises a discussion about the specific improvements our framework brings to baseline LLMs. While the framework is designed to bridge the knowledge gap between trial criteria and LLMs’ existing knowledge, our experiments indicate that LLMs can already expand and interpret the criteria in our datasets accurately. Instead, the primary benefit of knowledge augmentation lies in explicitly organizing relevant concepts and definitions in a well-structured format. By clearly outlining these key elements, the augmentation provides a type of breakdown of the subsequent reasoning steps, leading to improved overall performance.
However, despite the dataset limitations, the effectiveness of the augmentation module remains significant. One of its key strengths is its potential to interface with diverse data sources, particularly those that are not publicly available, such as paid subscription-based medical databases or proprietary clinical trial repositories. Furthermore, it plays a crucial role in integrating the most up-to-date medical knowledge, including newly developed treatments and evolving FDA guidelines. This capability ensures that the framework remains adaptable and relevant, even when dealing with rapidly changing medical landscapes where LLMs alone may lack the latest information. By systematically incorporating external knowledge, the augmentation module enhances the contextual relevance of trial matching, making it particularly valuable in real-world applications where timely and comprehensive knowledge retrieval is essential.
6. Conclusion
In this study, we introduced MAKAR, a scalable multi-agent framework designed to enhance patient-trial matching by combining knowledge augmentation with structured reasoning. By integrating an automated augmentation module with a reasoning module, MAKAR refines complex trial criteria and systematically evaluates patient eligibility. Our experimental results demonstrate that MAKAR outperforms existing benchmarks, including the state-of-the-art closed-source reasoning LLMs, and achieve 100% accuracy in a real-world offline test. And it also achieves up to an 8% improvement on complex criteria in a public dataset. Beyond performance gains, our findings emphasize the importance of explicitly structuring and expanding trial criteria to bridge the knowledge gap in AI-driven clinical reasoning.
It is worth noting that the design of MAKAR supports the future development of a human-AI collaboration system. For the trial-centric patient matching task, the workload of criterion-supervision performed by the critic agent is one-time per trial and relatively small and can be efficiently handled by human experts with minimal time investment. Exploring how humans can further enhance and facilitate this system presents an interesting direction for future research.
To address the limitations of this study and validate MAKAR’s generalizability, future work could also focus on evaluating the framework using more challenging datasets, such as those involving newly launched oncology trial that required up-to-date medical knowledge. Testing MAKAR in thess contexts would not only provide a more rigorous assessment of its capabilities but also demonstrate its potential to address critical real-world challenges in clinical trial recruitment.
7. Acknowledgments
We gratefully acknowledge support from the OpenAI Research Access Program, which provided access to advanced language models that were instrumental in developing and evaluating the framework presented in this work.
Appendix A Criteria
A.1. N2C2 Criteria Statistics
Criterion Met Count Not Met Count
ABDOMINAL 107 181
ADVANCED-CAD 170 118
ALCOHOL-ABUSE 10 278
ASP-FOR-MI 230 58
CREATININE 106 182
DIETSUPP-2MOS 149 139
DRUG-ABUSE 15 273
ENGLISH 265 23
HBA1C 102 186
KETO-1YR 1 287
MAJOR-DIABETES 156 132
MAKES-DECISIONS 277 11
MI-6MOS 26 262
Table 5. Counts of Eligible and Ineligible Patients for Each Criterion
A.2. N2C2 Original Criteria
Criteria Definition
DRUG-ABUSE Drug abuse, current or past.
ALCOHOL-ABUSE Current alcohol use over weekly recommended limits.
ENGLISH Patient must speak English.
MAKES-DECISIONS Patient must make their own medical decisions.
ABDOMINAL History of intra-abdominal surgery, small or large intestine resection, or small bowel obstruction.
MAJOR-DIABETES Major diabetes-related complications, such as: amputation, kidney damage, skin conditions, retinopathy, nephropathy, neuropathy.
ADVANCED-CAD Advanced cardiovascular disease, defined as two or more of the following: medications for CAD, history of myocardial infarction, angina, ischemia.
MI-6MOS Myocardial infarction in the past 6 months.
KETO-1YR Diagnosis of ketoacidosis in the past year.
DIETSUPP-2MOS Taken a dietary supplement (excluding Vitamin D) in the past 2 months.
ASP-FOR-MI Use of aspirin to prevent myocardial infarction.
HBA1C Any HbA1c value between 6.5 and 9.5%.
CREATININE Serum creatinine ¿ upper limit of normal.
Table 6. Definitions of the original criteria in the n2c2 dataset.
A.3. Redefined N2C2 Criteria
Table 7. Redefined Criteria proposed by (Wornow et al., 2024)
Criteria Definition
ABDOMINAL History of intra-abdominal surgery. This could include any form of intra-abdominal surgery, including but not limited to small/large intestine resection or small bowel obstruction.
ADVANCED-CAD Advanced cardiovascular disease (CAD). Defined as having 2 or more of the following: (a) Taking 2 or more medications to treat CAD (b) History of myocardial infarction (MI) (c) Currently experiencing angina (d) Ischemia, past or present. The patient must have at least 2 of these categories (a,b,c,d) to meet this criterion, otherwise the patient does not meet this criterion. For ADVANCED-CAD, be strict in your evaluation of the patient – if they just have cardiovascular disease, then they do not meet this criterion.
ALCOHOL-ABUSE Current alcohol use over weekly recommended limits.
ASP-FOR-MI Use of aspirin for preventing myocardial infarction (MI).
CREATININE Serum creatinine level above the upper normal limit.
DIETSUPP-2MOS Consumption of a dietary supplement (excluding vitamin D) in the past 2 months. To assess this criterion, go through the list of medications and supplements taken from the note. If a substance could potentially be used as a dietary supplement (i.e., it is commonly used as a dietary supplement, even if it is not explicitly stated as being used as a dietary supplement), then the patient meets this criterion. Be lenient and broad in what is considered a dietary supplement. For example, a ’multivitamin’ and ’calcium carbonate’ should always be considered a dietary supplement if they are included in this list.
DRUG-ABUSE Current or past history of drug abuse.
ENGLISH Patient speaks English. Assume that the patient speaks English, unless otherwise explicitly noted. If the patient’s language is not mentioned in the note, then assume they speak English and thus meet this criteria.
HBA1C Any hemoglobin A1c (HbA1c) value between 6.5% and 9.5%.
KETO-1YR Diagnosis of ketoacidosis within the past year.
MAJOR-DIABETES Major diabetes-related complication. Examples of “major complication” (as opposed to “minor complication”) include, but are not limited to, any of the following that are a result of (or strongly correlated with) uncontrolled diabetes: • Amputation • Kidney damage • Skin conditions • Retinopathy • Nephropathy • Neuropathy. Additionally, if multiple conditions together imply a severe case of diabetes, then count that as a major complication.
MAKES-DECISIONS Patient must make their own medical decisions. Assume that the patient makes their own medical decisions, unless otherwise explicitly noted. There is no information provided about the patient’s ability to make their own medical decisions, then assume they do make their own decisions and therefore meet this criteria.
MI-6MOS Myocardial infarction (MI) within the past 6 months.
A.4. MAKAR-Augemented N2C2 Criteria
Table 8. Redefined Criteria proposed by (Wornow et al., 2024)
Criteria Definition
ABDOMINAL Criteria: History of intra-abdominal surgery, including small/large intestine resection or small bowel obstruction.
Explanation: Intra-abdominal surgery involves operations on organs within the abdominal cavity (e.g., stomach, intestines). Small bowel obstruction refers to blockages in the small intestine that may require surgery.
ADVANCED-CAD Criteria: Advanced cardiovascular disease (CAD) defined by having two or more of the following: (a) Taking 2 or more medications for CAD, (b) History of myocardial infarction (MI), (c) Current angina, or (d) Ischemia (past or present).
Explanation: CAD affects heart function due to blocked arteries. Definitions include: MI (heart attack), angina (chest pain), ischemia (reduced blood flow), and specific medications like statins or beta-blockers. A diagnosis of CAD alone does not qualify as advanced.
ALCOHOL-ABUSE Criteria: Current alcohol use exceeding weekly recommended limits.
Explanation: Weekly limits: up to 14 standard drinks. A standard drink is  14g of pure alcohol (12oz beer, 5oz wine, or 1.5oz spirits).
ASP-FOR-MI Criteria: Use of aspirin to prevent myocardial infarction (MI).
Explanation: Aspirin reduces MI risk by preventing clots but requires medical advice to balance benefits (risk reduction) and potential side effects (e.g., bleeding).
CREATININE Criteria: Serum creatinine level above the upper normal limit.
Explanation: Creatinine reflects kidney function. Normal levels vary by age and gender (e.g.,  0.6-1.2 mg/dL for adult males). Elevated levels suggest potential kidney impairment.
DIETSUPP-2MOS Criteria: Consumption of a dietary supplement (excluding Vitamin D) in the past 2 months.
Explanation: Supplements include multivitamins, calcium carbonate, fish oil, or herbal supplements like ginseng. Excludes Vitamin D explicitly.
DRUG-ABUSE Criteria: Current or past drug abuse history.
Explanation: Drug abuse includes misuse of psychoactive substances, with potential consequences like health issues or legal problems.
ENGLISH Criteria: Patient must speak English.
Explanation: English proficiency ensures effective participation in English-based trials. Assume English proficiency unless noted otherwise.
HBA1C Criteria: Hemoglobin A1c (HbA1c) between 6.5% and 9.5%.
Explanation: HbA1c measures average blood sugar over 2–3 months. The range identifies diabetes patients with moderate control, excluding poorly controlled cases.
KETO-1YR Criteria: Diagnosis of ketoacidosis in the past year.
Explanation: Ketoacidosis is a serious condition with high blood ketones, often related to uncontrolled diabetes. Symptoms include nausea, confusion, and rapid breathing.
MAJOR-DIABETES Criteria: Major diabetes-related complications like amputation, kidney damage, or neuropathy.
Explanation: Includes significant health effects (e.g., nephropathy, retinopathy) stemming from uncontrolled diabetes. Multiple minor complications combined may also qualify.
MAKES-DECISIONS Criteria: Patient must make their own medical decisions.
Explanation: This assesses cognitive ability to understand medical options, communicate choices, and appreciate consequences.
MI-6MOS Criteria: Myocardial infarction (MI) within the past 6 months.
Explanation: MI refers to heart muscle damage from blocked blood flow. Assess via medical history and records confirming the event within 6 months.
A.5. Sampled ClinicalTrial.gov Criteria
Inclusion Criteria Exclusion Criteria
1. Locally advanced or metastatic breast cancer confirmed by histopathology. 1. Patients known or suspected intolerance or hypersensitivity to main ingredient or any of the excipients of SNB-101.
2. Received 2 previous lines of anti-tumor treatment and developed resistance to standard treatment. 2. Patients with bilateral invasive breast cancers.
3. Patients with locally advanced or metastatic advanced solid tumors confirmed by histology or cytology (including but not limited to triple-negative breast cancer, gastric cancer, colorectal cancer); Patients with 1 line of standard treatment failure (disease progression after treatment or intolerability of toxic side effects of treatment), or no standard treatment, or unable to receive standard treatment. 3. Age $<$18 years old or $>$70 years old.
4. At least 1 measurable lesion per Response Evaluation Criteria in Solid Tumors version 1.1 and previously treated lesions with radiotherapy or focal therapy and no progression cannot be included as target lesion for assessment. 4. Patients with standard metallic contraindications to CMR or estimated glomerular filtration rate $<$ 30 mL/min/1.73 m².
5. Histologically confirmed breast cancer with an invasive component measuring 20 mm and/or with morphologically confirmed spread to regional lymph nodes (stage cT2-cT4 with any cN, or cN1-cN3 with any cT). 5. History of hypersensitivity or contraindication to TTF.
6. Known HER2-positive breast cancer defined as an IHC status of 3+. If IHC is 2+, a positive in situ hybridization (FISH, CISH, or SISH) test is required by local laboratory testing. 6. Implanted pacemaker, defibrillator, or other electrical medical devices.
7. Patients with breast cancer planned to receive Anthracycline and Cyclophosphamide chemotherapy. 7. Bilateral breast cancer (including multifocal breast cancer. All histological lesions were HER2 1+ or 2+, not detected by FISH amplification and HR positive according to ASCO guidelines).
8. Patients treated for hormone-dependent localized breast cancer requiring adjuvant hormonal therapy (HT). 8. Received any form of anti-tumor therapy (chemotherapy, radiotherapy, molecular targeted therapy, endocrine therapy, etc.).
9. Patients treated with tamoxifen for a maximum of 1 to 3 years. 9. Breast cancer without histopathological diagnosis.
10. Histologically confirmed ER+/HER2- early-stage resected invasive breast cancer at high or intermediate risk of recurrence, , based on clinical-pathological risk features, as defined in the protocol. 10. Known personal history of ductal carcinoma in situ (DCIS) or invasive breast cancer.
11. Completed adequate (definitive) locoregional therapy (surgery with or without radiotherapy) for the primary breast tumour(s), with or without (neo)adjuvant chemotherapy. 11. Prior systemic treatment for any malignancy.
12. Completed at least 2 years but no more than 5 years (+3 months) of adjuvant ET (+/- CDK4/6 inhibitor). 12. Significant medical comorbidities as per investigator evaluation.
13. Patients receiving letrozole for more than two months. 13. Patients with ongoing indications for the cardioprotective medication - ACE inhibitors, ARBs and/or beta-blockers
14. Postmenopausal hormone receptor-positive patients. 14. Intermediate or high-grade ductal carcinoma in situ.
15. Patients with N0 disease or those who have undergone targeted axillary detection with a good response. 15. Invasive carcinoma.
Table 9. Sampled Criteria from Breast Cancer Clinical Trials

### v3

@@
-Enhancing Clinical Trial Patient Matching through Knowledge Augmentation and Reasoning with Multi-Agents
-Abstract.
+Enhancing Clinical Trial Patient Matching through Knowledge Augmentation and Reasoning with Multi-Agent
+Abstract
 Matching patients effectively and efficiently for clinical trials is a significant challenge due to the complexity and variability of patient profiles and trial criteria.
-This paper introduces Multi-Agents for Knowledge Augmentation and Reasoning (MAKAR), a novel framework that enhances patient-trial matching by integrating domain-specific knowledge with structured reasoning.
-MAKAR consists of two key modules: the Augmentation Module and the Reasoning Module.
-The Augmentation Module enriches trial criteria by incorporating detailed explanations of relevant concepts to ensure clarity and completeness.
-The Reasoning Module then evaluates each health condition, following a structured, step-wise approach to determine eligibility and make the final matching decision.
-This paper enhances patient-trial matching by leveraging the agency and reasoning capabilities of large language models (LLMs) through automated agent interactions, including collaboration, critique, and navigation.
-Experimental results on a public dataset demonstrate that our framework surpasses existing benchmarks, achieving up to an 8% improvement in accuracy for specific criteria.
-Furthermore, in a real-world offline test, MAKAR achieved a 100% accuracy.
-These findings show MAKAR’s potential as a scalable and robust solution for clinical trial patient matching.
-Patient Matching, Multi-agent Systems, Knowledge Augmentation, Reasoning LLMs
-ccs: Computing methodologies Multi-agent systemsccs: Applied computing Health care information systems
-1. Introduction
-Patient matching for clinical trials refers to the process of identifying and enrolling participants whose health profiles meet the specific eligibility criteria of a given clinical study.
+This paper introduces Multi-Agent for Knowledge Augmentation and Reasoning (MAKAR), a novel multi-agent system that enhances patient-trial matching by integrating criterion augmentation with structured reasoning.
+MAKAR consistently improves performance by an average of 7% across different datasets.
+Furthermore, it enables privacy-preserving deployment and maintains competitive performance when using smaller open-source models.
+Overall, MAKAR can contributes to more transparent, accurate, and privacy-conscious AI-driven patient matching.
+Enhancing Clinical Trial Patient Matching through Knowledge Augmentation and Reasoning with Multi-Agent
+Hanwen Shi University of Maryland hwshi@umd.edu                        Jin Zhang Pi Health Jin.Zhang@pihealth.ai                        Kunpeng Zhang University of Maryland kpzhang@umd.edu
+Figure 1: Comparison of accuracy of zero-shot patient matching with different prompts. The values are averaged over 5 runs; error bars show the observed range.
+1 Introduction
+Patient matching for clinical trials is the process of identifying and enrolling participants whose health profiles meet the specific eligibility criteria of a given clinical study.
 Efficient patient matching is crucial to accelerating the drug development process and reducing the overall cost of clinical trials.
@@
 The emergence of of LLMs offers a promising opportunity to improve patient matching.
-LLMs, such as GPT-4 (Achiam et al., 2023), usually possess a large text window and advanced natural language understanding capabilities (Karanikolas et al., 2023), which enable them process complex trial criteria and patient data effectively.
-However, their general knowledge may not always be sufficient for knowledge-intensive tasks like patient matching in clinical trials, where domain-specific expertise is required.
-The knowledge gap inherent in LLMs and misssing information in the pre-defined criteria can limit their performance, highlighting the need for LLM augmentation of specialized medical knowledge (Shuster et al., 2021).
-While research indicates that LLMs do encode medical knowledge in a parameterized form (Singhal et al., 2023), this information can be outdated or misaligned with the special needs of certain clinical trials.
-Figure 1. The workflow of MAKAR framework
-To address these limitations, we present a new framework called: Multi-Agents for Knowledge Augmentation and Reasoning (MAKAR).
-MAKAR is a novel framework designed to enhance the patient matching process.
-The framework consists of two core modules: the Augmentation Module and the Reasoning Module.
-The Augmentation Module enhances trial criteria by providing detailed explanations and includes three key components: a Router Agent, multiple Augmentation Agents, and a Critic Agent.
-The Reasoning Module systematically evaluates each criterion-related condition in a step-by-step manner to determine eligibility and make the final matching decision.
-It comprises a Reasoning Agent and a Supervision Matching Agent.
-By leveraging this multi-agent framework, MAKAR improves the performance on patient-trial matching by dynamically incorporating domain-specific knowledge and step-wise reasoning.
-Our experimental results demonstrate the effectiveness of MAKAR.
-On a public dataset, our framework achieves an average improvement of 3% over benchmark methods, with up to an 8% improvement for complex criteria.
-Furthermore, in an offline real-world data test, MAKAR outperformed the state-of-the-art closed-source reasoning LLMs and achieved 100% accuracy, showing its reliability in practical applications.
+LLMs, such as GPT-4 Achiam et al.
+(2023), usually possess a large text window and advanced natural language understanding capabilities (Karanikolas et al., 2023), which enable them process complex trial criteria and patient data effectively.
+However, three limitations constrain the performance of LLMs.
+First, knowledge gaps (Shuster et al., 2021) may exist due to training data cut-off dates and incomplete coverage of specialized medical knowledge, such as newly approved drugs, treatments, or updated FDA guidelines.
+Second, hallucinations (Huang et al., 2025) remain a significant challenge: LLMs are often prompted to generate detailed reasoning to improve interpretability and trustworthiness in patient matching (Jin et al., 2024), yet such reasoning is unverified and potentially misleading.
+Finally, prompt sensitivity can substantially impact outputs, as even minor variations in prompts can lead to divergent results (Zhuo et al., 2024).
+To address these limitations, we present a training-free multi-agent workflow: Multi-Agent for Knowledge Augmentation and Reasoning (MAKAR).
+MAKAR consists of two multi-agent modules: the Augmentation Module and the Reasoning Module.
+The augmentation module enriches trial criteria by providing detailed explanations from different sources.
+And the reasoning module evaluates each criterion-related patient condition in a step-by-step manner to determine eligibility and make the final matching decision.
+We demonstrate the effectiveness of MAKAR our framework on public dataset and in-house real-world database.
+On average, MAKAR achieves an improvement of 7% over benchmark methods, with up to a 10% improvement on certain criteria.
 Our key contributions are as follows:
-  • We propose an automated method to refine and enhance trial criteria, making them more interpretable and suitable for AI-driven patient-trial matching.
-  • We introduce a multi-agent framework that combines augmentation and reasoning, achieving strong performance across both public and real-world datasets.
-  • We identify a gap in existing approaches and highlight the importance of explicitly incorporating well-structured, detailed explanations of related concepts to improve LLM-driven clinical reasoning.
-2. Related Work
-2.1. Clinical Trial Patient Matching
-Traditional patient matching systems rely on query-based structured databases, such as Criteria2Query (Yuan et al., 2019), or handcrafted rule-based approaches like RBC (Oleynik et al., 2019).
-However, these methods struggle to handle unstructured free-text data and are highly sensitive to variations in formatting and verbal expressions.
-Unfortunately, a significant portion of EHRs consists of free-text, including clinical notes, surgical records, radiology reports, pathology reports and so on (Sun et al., 2018).
-Thus, the effectiveness and scalability of traditional methods are significantly limited, making them unsuitable for future patient matching tasks.
-Recent advancements in LLMs have shown remarkable capabilities in directly processing free-text clinical trial criteria, offering scalable solutions without the need for extensive rule-based preprocessing.
-Many Studies have applied these models to patient-centric trial matching tasks using real-world data.
-For instance, Nievas et al.
-(2024) fine-tuned the open-source Llama 2 (Touvron et al., 2023) to automate patient matching.
-However, fine-tuning LLMs is computationally expensive, prompting alternative approaches that leverage frozen models.
-Jin et al.
-(2023) introduced TrialGPT, a patient-centric matching framework that assesses patient eligibility on a criterion-by-criterion basis, filtering out ineligible trials and ranking potentially eligible ones.
-Their results indicated a strong correlation between TrialGPT’s predictions and expert clinician judgments, achieving accuracy at the criterion level similar to human experts.
-Similarly, Gupta et al.
-(2024) proposed OncoLLM, an end-to-end pipeline that converts trial criteria into conversational prompts and uses a Chain-of-Thought (CoT) (Wei et al., 2022) reasoning strategy.
-OncoLLM outperforms GPT-3.5 while underperforms GPT-4, despite its smaller model size.
-In contrast, other studies have focused on trial-centric patient matching, typically using public deidentified datasets due to HIPAA-related data constraints.
-Wornow et al.
-(2024) developed a LLM-based system with a two-stage retrieval pipeline to identify suitable patients for trials using a zero-shot prompt approach.
-It reduces token processing costs while outperforming existing state-of-the-art models on the n2c2 2018 cohort selection benchmark (Stubbs et al., 2019).
-Likewise, Beattie et al.
-(2024) applied a CoT strategy in a Retrieval-Augmented Generation (RAG) (Lewis et al., 2020) system using GPT-3.5 Turbo and GPT-4, demonstrating results comparable to those of Wornow et al.
-(2024).
-These works collectively highlight the potential of LLMs in clinical trial matching.
-However, existing approaches directly employ LLMs without verifying whether the trial criteria are sufficiently understood by the model (Webson and Pavlick, 2021) or whether there are gaps in the model’s knowledge (Lewis et al., 2020).
-Our work fits within the trial-centric patient matching literature and makes a novel contribution by automatically identifying and addressing these gaps, thereby enhancing the accuracy patient matching.
-2.2. LLM-based Multi-Agents (LLM-MA)
-LLM-MA have seen rapid development across various domains, driven by the promising potential of LLMs as autonomous agents.
-Guo et al.
-(2024) provide a comprehensive overview of recent advancements in LLM-MA applications, categorizing them into two primary areas: Problem-Solving and World Simulation.
-Problem-solving LLM-MA systems explore the capabilities of agents with specialized expertise, enabling them to outperform single-agent LLM systems through collaborative efforts.
-For instance,  Wang et al.
-(2024) proposed Mixture-of-Agents, a layered multi-agent architecture that achieves state-of-the-art performance in benchmarks like AlpacaEval 2.0 (Dubois et al., 2024).
-In the healthcare domain, the application of LLM-MA systems is rapidly gaining traction.
-Yue et al.(Yue and Fu, 2024) introduce ClinicalAgent, a multi-agent architecture powered by LLM reasoning for multi-agent collaboration, achieving higher precision in predicting clinical trial outcomes.
-Similarly, Liao et al.(Liao et al., 2024) propose REFLECTool, a multi-agent framework that enables agents to learn how to select domain-specific tools, thereby extend their capabilities.
-TriageAgent(Lu et al., 2024) leverages LLMs for role-playing interactions, integrating self-confidence assessments and early-stopping mechanisms in multi-round discussions to improve document reasoning and classification accuracy for triage tasks.
-Tang et al.(Tang et al., 2023) use multiple LLM-based agents as domain experts to collaboratively analyze medical reports.
-And Li et al.
-(Li et al., 2024) further introduce the multi-agent collaboration framework to handle multimodal tasks.
-To the best of our knowledge, this study is the first to apply an LLM-MA framework for cohort selection in clinical trials.
-Our approach explore the agency and reasoning capabilities of LLM-MA to dynamically enhance patient-trial matching with tailored knowledge.
-Our approach allows a better contextual understanding of patient profiles and lead to improved matching performance and more informed decision-making in clinical trial recruitment.
-Criteria Accuracy Precision Recall F1-score
-Zero-shot CoT MAKAR Zero-shot CoT MAKAR Zero-shot CoT MAKAR Zero-shot CoT MAKAR
-ABDOMINAL 0.858 0.868 0.847 0.893 0.897 0.944 0.701 0.729 0.626 0.785 0.804 0.753
-ADVANCED-CAD 0.854 0.854 0.868 0.895 0.872 0.929 0.853 0.882 0.841 0.873 0.877 0.883
-ALCOHOL-ABUSE 0.969 0.969 0.976 0.529 0.545 0.667 0.900 0.600 0.600 0.667 0.571 0.632
-ASP-FOR-MI 0.861 0.875 0.885 0.899 0.908 0.919 0.930 0.939 0.939 0.915 0.923 0.929
-CREATININE 0.802 0.802 0.885 0.664 0.667 0.812 0.934 0.925 0.896 0.776 0.775 0.852
-DIETSUPP-2MOS 0.764 0.781 0.785 0.704 0.717 0.721 0.940 0.953 0.953 0.805 0.818 0.821
-DRUG-ABUSE 0.941 0.920 0.993 0.469 0.395 0.882 1.000 1.000 1.000 0.638 0.566 0.938
-ENGLISH 0.976 0.986 0.983 0.985 0.989 0.996 0.989 0.996 0.985 0.987 0.992 0.991
-HBA1C 0.840 0.830 0.878 0.722 0.702 0.807 0.892 0.902 0.863 0.798 0.790 0.834
-KETO-1YR 0.993 0.990 0.993 - - - - - - - - -
-MAJOR-DIABETES 0.833 0.826 0.847 0.851 0.831 0.873 0.840 0.853 0.840 0.845 0.842 0.856
-MAKES-DECISIONS 0.920 0.924 0.931 0.985 0.985 0.996 0.931 0.935 0.931 0.957 0.959 0.963
-MI-6MOS 0.885 0.872 0.927 0.434 0.396 0.571 0.885 0.808 0.769 0.582 0.532 0.656
-Average 0.884 0.884 0.908 0.750 0.740 0.841 0.899 0.876 0.853 0.801 0.786 0.841
-Table 1. Criterion-Level Comparison Across Different Methods for N2C2
-3. Methods
-The proposed framework, shown in Figure 1, consists of two core modules: the Augmentation Module and the Reasoning Module.
-The augmentation module is designed to generate criteria with comprehensive explanations of all relevant concepts, while the reasoning module evaluates each condition within the criteria to determine eligibility and make the final matching decision.
-Additionally, the framework contains five interconnected components, each outlined below.
-Router Agent In the MAKAR framework, multiple specialized augmentation agents are available to augment the criterion with domain-specific knowledge.
+  • We demonstrate by experiment that LLMs are sensitive to how criteria are formulated.
+  • We introduce MAKAR, a double-module multi-agent workflow that allows complementary contributions from criterion augmentation and reasoning.
+And it achieves consistent performance improvement across different datasets.
+  • We demonstrate that MAKAR remains effective when deployed with open-source language models.
+And it has potential to contribute to more trustworthy and privacy-preserving AI application in drug development.
+2 Criterion Sensitivity
+To further motivate this work and demonstrate existing gaps, we conducted a preliminary experiment comparing patient matching performance on the 2018 n2c2 Challenge Track 1 dataset (Stubbs et al., 2019) using three different criteria: (1) the original criteria provided by the challenge (see Appendix B.1), (2) human-refined criteria used by Wornow et al.
+(2024) (Appendix B.2), and (3) GPT-4o-refined criteria (Appendix B.3).
+All experiments were performed with GPT-4o in a zero-shot setting, and the prompts are provided in Appendix C.1.
+The prompt used to refine the criteria in (3) is provided in Appendix C.2.
+The results are presented in Figure 1.
+Overall, the results in Figure 1 suggests three important observations.
+First, the criterion sensitivity of LLM-based patient matching is substantial: simply varying the formulation of the criteria can cause notable differences in accuracy across most conditions.
+Second, while using human-refined criteria does not universally outperform the original criteria, automatically refined and augmented prompts generated by GPT-4o consistently improve performance.
+Third, explicitly incorporating detailed knowledge (e.g., explanations) into the prompt can further enhance accuracy, even when the LLMs themselves do not exhibit an obvious knowledge gap.
+For example, for the "CREATININE" criterion, GPT-4o was able to return the correct reference ranges by gender and age when being asked directly.
+However, explicitly adding these ranges in the prompt—as done in the GPT-4o-augmented criteria—increased performance by 8% compared to prompts that used only general descriptions such as “normal range.” These findings highlight the need for methods that can systematically augment knowledge rather than relying solely on prompt engineering or implicit model knowledge.
+Figure 2: The workflow of MAKAR framework
+3 Methods
+Our proposed methods, shown in Figure 2, is a training-free multi-agent workflow and consists of two core modules: the Augmentation Module and the Reasoning Module.
+The augmentation module is designed to generate criteria with comprehensive explanations of relevant concepts, while the reasoning module evaluates each condition within the criteria to determine eligibility and make the final matching decision.
+The workflow contains five interconnected components and all prompts used by the agents listed below can be found in Appendix C.3.
+The pseudo code of the workflow is provided in Appendix D
+Router Agent In MAKAR, multiple specialized augmentation agents are available to augment the criterion with domain-specific knowledge.
 The Router Agent determines the most suitable augmentation agent for the task and routes the workflow accordingly.
-Inspired by SELFREF (Kadavath et al., 2022) and MOREINFO (Feng et al., 2023), we adopt a prompt-based self-probing strategy.
+Inspired by SELFREF Kadavath et al.
+(2022) and MOREINFO Feng et al.
+(2023), we adopt a prompt-based self-probing strategy.
 If the agent determines that the original criterion is detailed and understandable enough for other agents, the criterion will be directly forwarded to the Reasoning Module.
@@
 Knowledge Augmentation Agent: This component augments the criterion with relevant, domain-specific information from different source.
-For instance, the Retrieval Agent (Lewis et al., 2020) retrieves information from indexed medical databases, the Online Search Agent gathers supplementary knowledge from online sources, and the Self-Augment Agent leverages the LLM’s intrinsic knowledge to refine the original criterion.
-Critic Agent: Acting as a supervisor, this agent ensures that the augmented criterion is aligned with the original instructions.
+For instance, the Retrieval Agent Lewis et al.
+(2020) retrieves information from indexed medical databases, the Online Search Agent gathers supplementary knowledge from online sources, and the Self-Augment Agent leverages the LLM’s intrinsic knowledge to refine the original criterion.
+Critic Agent: We add Critic Agent as the supervisor to power the iterative refinement with self-feedback as in SELF-REFINE Madaan et al.
+(2023).
+It ensures that the augmented criterion is aligned with the original instructions.
 For instance, the original criterion must be preserved verbatim.
@@
 Reasoning Agent With the augmented criterion and patient EHR data, the Reasoning Agent evaluates patient eligibility by assessing for each concepts mentioned in the augmented criterion.
-It generates a detailed reasoning summary, including assessments for all relevant health conditions, and forwards it to the Supervision Matching Agent.
+It generates a detailed reasoning summary Jin et al.
+(2024), including assessments for all relevant health conditions, and forwards it to the Matching Agent.
 Matching Agent: The Matching Agent determines patient eligibility by reviewing the reasoning assessment alongside the patient’s EHR data, making a final judgment in a zero-shot manner.
-4. Evaluation & Results
-Due to the sensitive nature of patient data, which is strictly protected under The Health Insurance Portability and Accountability Act (HIPAA), and the compliance requirements for processing detail clinical trial documents, full-scale online deployment is not currently feasible—particularly due to concerns about data security when using closed-source LLMs.
-Instead, to evaluate the effectiveness of our methods, we conducted two experiments: (1) an assessment using a public dataset and (2) a limited-scale offline evaluation with deidentified real-world data.
-4.1. Data
-The public dataset we use is from Track 1 of the 2018 n2c2 challenge (Stubbs et al., 2019), which is commonly used for evaluating patient matching in clinical trials.
+Criteria Jin et al. (2024) Wornow et al. (2024)  Beattie et al. (2024) MAKAR
+Accuracy F1 Accuracy F1 Accuracy F1 Accuracy F1
+N2C2 Dataset
+ABDOMINAL $0.813$ $0.722$ $0.858$ $0.785$ 0.868 0.804 $0.809$ $0.671$
+ADVANCED-CAD $0.823$ $0.859$ $0.854$ $0.873$ $0.854$ $0.877$ 0.882 0.897
+ALCOHOL-ABUSE $0.972$ $0.667$ $0.969$ $0.667$ $0.969$ $0.571$ 0.983 0.706
+ASP-FOR-MI 0.906 0.944 $0.861$ $0.915$ $0.875$ $0.923$ $0.896$ $0.938$
+CREATININE $0.802$ $0.776$ $0.802$ $0.776$ $0.802$ $0.776$ 0.924 0.900
+DIETSUPP-2MOS $0.722$ $0.759$ $0.764$ $0.805$ $0.781$ $0.818$ 0.826 0.846
+DRUG-ABUSE $0.882$ $0.469$ $0.941$ $0.638$ $0.920$ $0.566$ 0.972 0.765
+ENGLISH $0.948$ $0.971$ $0.976$ $0.987$ $0.986$ $0.992$ 0.997 0.998
+HBA1C $0.840$ $0.800$ $0.840$ $0.798$ $0.830$ $0.790$ 0.934 0.899
+KETO-1YR 0.993 - 0.993 - $0.990$ 0.993 -
+MAJOR-DIABETES $0.830$ $0.838$ $0.833$ 0.845 $0.826$ $0.842$ 0.840 $0.838$
+MAKES-DECISIONS $0.580$ $0.722$ $0.920$ $0.957$ $0.924$ $0.959$ 0.965 0.982
+MI-6MOS $0.833$ $0.510$ $0.885$ $0.582$ $0.872$ $0.532$ 0.962 0.807
+Average $0.842$ $0.753$ $0.884$ $0.802$ $0.884$ $0.787$ 0.922 0.854
+ClinicalTrial Dataset
+CT-IN $0.900$ $0.828$ $0.967$ $0.929$ 0.987 0.982
+CT-EX $0.900$ $0.885$ $0.947$ $0.937$ 0.987 0.982
+Table 1: Performance Comparison Across Models on N2C2 and ClinicalTrial Datasets.
+CT-IN and CT-EX stand for inclusion and exclusion criteria in the ClinicalTrial dataset.
+The F1 score for KETO-1YR is missing because only one patient met this criterion, and no method match it correctly.
+4 Experiments
+4.1 Data
+We use a widely used public dataset and an in-house real-world dataset to evaluate our proposed method.
+Both datasets were labeled by at least two trained annotators, with any disagreements resolved by an independent physician to determine the final ground truth.
+The public dataset is from Track 1 of the 2018 n2c2 challenge Stubbs et al.
+(2019).
 The dataset consists of 288 de-identified patients, each having between 2 to 5 unstructured longitudinal clinical notes in American English.
-The entire corpus contains 781,006 tokens, with an average of 2,711 tokens per patient.
-The challenge simulates a synthetic clinical trial with 13 predefined inclusion criteria, resulting in 3744 patient-criteria pairs in total.
-The detailed criteria can be found in Table 6.
-For each patient-criteria pair, a binary label is provided, indicating whether the patient ”MET” or ”NOT MET” the criterion.
-Two annotators with medical background labeled these data.
-The disagreements between the two annotators were resolved by an independent MD to establish the final ground truth labels.
-The criteria in the n2c2 dataset are relatively straightforward and the patient records are well-structured, making them easy to follow.
-To evaluate the potential of our proposed method in a more complex scenario, we constructed a real-world dataset.
-Specifically, we sampled 30 eligibility criteria—15 inclusion and 15 exclusion—from ongoing breast cancer trials listed on ClinicalTrials.gov.
-We then manually selected 10 patients who have been diagnosed with breast cancer from our in-house EHR system.
-For each patient, we extracted their most recent deidentified clinical visit record.
-An expert subsequently asseessed whether each patient met each criterion, resulting in 300 labeled criterion-patient pairs.
-Notably, if a patient’s information was insufficient to determine eligibility for a given criterion, they were labeled as ”NOT MET” to ensure a conservative patient selection.
-We listed the sampled criterion in Table  9
-4.2. Baselines and Metrics
-Previous research has demonstrated the superiority of LLM-based methods over traditional rule-based systems; hence, this study focuses exclusively on LLM-based benchmarks.
-We include two approaches in Experiment 1: the zero-shot method with hand-crafted criteria (Wornow et al., 2024), as shown in Table LABEL:table:redefined_criteria, and a Chain-of-Thought (CoT) approach (Beattie et al., 2024) using the original criteria.
-In addtion, we also include o3-mini from OpenAI as the benchmark in Experiment 2 since reasoning LLMs have shown significant improvement on complex clinical scenarios(Xie et al., 2024).
-In line with prior literature, we report accuracy, precision, recall, and F1-score for this binary classification task.
-Previous studies using n2c2 dataset have primarily focused on evaluating classification performance at the individual criterion level, often overlooking the overall patient-trial assessment.
-In this study, we also examine patient-trial selection performance in Experiment 1.
-Specifically, a patient is eligible for the trial only if all inclusion criteria are met.
-However, no patient in the n2c2 dataset satisfied all 13 criteria.
-To mitigate this, we synthesize a trial with selected criteria that had at least 100 patients meeting them, resulting in 28 eligible patients from a total cohort of 288.
-We then evaluated our method’s on this new synthetic trial in identifying eligible patients and reported the findings.
-4.3. Experiment Settings
-Our experiments were conducted on an Amazon EC2 g5.16xlarge instance, equipped with a A10G GPU with 24 GB of memory. Unless otherwise specified, all agents in the framework are powered by GPT-4o.
+The challenge simulates a synthetic clinical trial with 13 predefined inclusion criteria (see Appendix B.1), resulting in 3744 patient-criteria pairs.
+Each pair has a binary label—“MET” or “NOT MET” and the statistics can be found in Table 4.
+To assess our method in a more complex setting, we created a real-world dataset by sampling 30 eligibility criteria (15 inclusion, 15 exclusion) from ongoing breast cancer trials on ClinicalTrials.gov.
+We manually selected 10 breast cancer patients from our in-house EHR system and extracted their most recent deidentified clinical records.
+When information was insufficient to confirm eligibility, the criterion was conservatively labeled as "NOT MET." The sampled criteria are listed in Appendix B.4.
+4.2 Baselines and Metrics
+Previous research has shown that LLM-based approaches outperform traditional rule-based systems Jin et al.
+(2024).
+Thus, this study focuses exclusively on LLM-based benchmarks.
+We include three benchmarks: (1) TrialGPT Jin et al.
+(2024), (2) the two-stage retrieval-based zero-shot matching system from Wornow et al.
+(2024), and (3) the chain-of-thought (CoT) based method introduced by Beattie et al.
+(2024).
+For the n2c2 dataset, we report accuracy and F1-score for each criterion as well as the overall average.
+For the ClinicalTrial dataset, we report the average performance for inclusion and exclusion criteria separately.
+4.3 Experiment Settings
+Our experiments were conducted on 4 A10G GPU with 96 GB of memory. Unless otherwise specified, all agents in the framework are powered by GPT-4o.
 For external knowledge augmentation, the Online Search Agent uses the Perplexity AI search engine to fetch up-to-date information from the web.
-The Retrieval Agent uses all-MiniLM-L12-v2, a pre-trained checkpoint from Sentence-Transformer (Reimers, 2019), as its encoder to construct the vector database of MedDRA and FDALabel.
-Metric Zero-shot CoT MAKAR
-Accuracy 0.8924 0.9201 0.9306
-Precision 0.4615 0.5676 0.6333
-Recall 0.6429 0.7500 0.6786
-F1-Score 0.5373 0.6462 0.6552
-Table 2. Trial-Level Comparison Across Different Methods
-4.4. Results of Experiment 1
-4.4.1. Criterion-Level Performance:
-Table 1 shows the performance of three approaches—Zero-shot,CoT, MAKAR—across multiple individual criteria.
-Our results indicate that while the CoT method slightly outperforms Zero-shot in Accuracy and Recall, MAKAR consistently achieves higher scores, particularly in Accuracy (0.909) and Precision(0.822) on average.
-The F1-score of MAKAR (0.828) suggests a balanced performance across criteria, although it falls slightly behind in a few specific cases such as ”ABDOMINAL” and ”English”.
-The MAKAR-augmented criteria, presented in Table LABEL:tab:aug_criteria, demonstrate significant improvements over the baseline human-redefined criteria outlined in Table LABEL:table:redefined_criteria from Wornow et al.
+The Retrieval Agent uses all-MiniLM-L12-v2, a pre-trained checkpoint from Sentence-Transformer Reimers (2019), as its encoder to construct the vector database of MedDRA and FDALabel.
+5 Results
+Table 1 compares the performance of MAKAR with benchmark methods across two datasets.
+Overall, the MAKAR model consistently achieves the highest or near-highest accuracy and F1-scores across most conditions.
+MAKAR improves accuracy by approximately 4% and F1-score by around 7% on average.
+Notably, it shows particularly strong gains on criteria such as CREATININE, HBA1C, and MI-6MOS, achieving improvements of about 10% where other methods struggle.
+To better understand these gains, we examine the MAKAR-augmented criteria presented in Appendix B.3.
+Compared to the baseline human-refined criteria outlined in Appendix B.2, MAKAR has two major improvements: (1) more detailed expansions of key concepts, for example, supplementing normal ranges for criteria like HBA1C and CREATININE across diverse demographic groups; and (2) more comprehensive and accessible definitions, such as using widely recognized terms like “heart attack” alongside the clinical term “myocardial infarction (MI)”.
+These improvements improve the ability of LLMs to match patients from clinical notes written in various styles, resulting in substantial performance gains on these criteria.
+Patient matching is not merely a binary classification task; the correctness of the reasoning process is essential for transparency, which in turn affects physicians’ trust in AI systems Asan et al.
 (2020).
-These enhancements primarily focus on providing explicit clarifications and detailed expansions of key concepts, enabling LLMs to better align with trial requirements.
-For example, MAKAR supplements normal ranges for criteria like ”HBA1C” and ”CREATININE” across diverse demographic groups.
-And they are also the criteria that benefit the most from MAKAR’s augmentation.
-Additionally, MAKAR provides a more comprehensive definition for criteria like ”MAJOR-DIABETES,” and use more widely recognized terminology, ”heart attack”, to explain clinical terms ”Myocardial Infarction (MI)”.
-These refinements improve the ability of LLMs to match patients from clinical notes written in a variety of styles, resulting in notable performance gains for these criteria.
-4.4.2. Trial-Level Performance
-The result of trial-levle analysis can be found in Table 2.
-At the trial level, where a patient’s eligibility is determined by satisfying all selected criteria, MAKAR outperforms both Zero-shot and CoT.
-MAKAR achieves the highest Accuracy (0.9306), Precision (0.6333)and F1-score (0.6552) but underperform CoT on Recall.
-The trial-level results further validate MAKAR’s capability in maintaining a higher balance between identifying eligible patients and minimizing patient-filtering load
-Inclusion Criteria
-Metric GPT-4o o1 o3-mini CoT MAKAR
-Accuracy 0.9000 0.9333 0.8533 0.9667 1.000
-Precision 0.7884 0.8407 0.7211 0.9385 1.000
-Recall 0.9020 0.9413 0.8151 0.9208 1.000
-F1-Score 0.8281 0.8800 0.7516 0.9294 1.000
-Exclusion Criteria
-Metric GPT-4o o1 o3-mini CoT MAKARCorr
-Accuracy 0.9000 0.9267 0.9133 0.9467 1.000
-Precision 0.8693 0.8987 0.8840 0.9265 1.000
-Recall 0.9090 0.9416 0.9253 0.9487 1.000
-F1-Score 0.8845 0.9153 0.8999 0.9365 1.000
-Table 3. Criterion-Level Comparison Across Different Methods in Experment 2
-4.5. Results of Experiment 2
-4.5.1. Results
-Table 3 presents the evaluation results of our method on the in-house real-world dataset.
-We compared its performance against zero-shot prompting with GPT-4o, zero-shot prompting with two reasoning models (OpenAI-o3-mini and OpenAI-o1) and CoT strategy with GPT-4o.
-Our method achieves perfect scores across all four metrics.
-Notably, for exclusion criteria, we report results for a corrected version of MAKAR (MAKARCorr).
-In the initial version, the model was not explicitly instructed on whether a given criterion was for inclusion or exclusion.
-Surprisingly, it was able to infer some exclusion criteria but focused on determining a patient’s overall eligibility for a trial rather than assessing whether a specific criterion was met.
-For instance, if a patient was 75 years old, they should have been marked as ”MET” for an exclusion criterion of ”greater than 70 years old”.
-While MAKAR correctly identified the age range and provided a logical reasoning process, it mistakenly output ”Not MET”—interpreting the task as determining overall trial eligibility rather than verifying the criterion itself.
-The corrected version addresses this by explicitly instructing the model to evaluate each criterion independently rather than making a broader eligibility judgment.
-MAKAR outperforms all the benchmarks and the CoT follows closely, achieving 96.67% accuracy and 92.94% F1-score for inclusion criteria and 94.67% accuracy with 93.65% F1-score for exclusion criteria.
-o3-mini, a smaller model, is outperformed by o1 but shows competitive performance on exclusion criteria with zero-shot method with GPT-4o.
-Interestingly, Experiment 2 yields better overall performance than Experiment 1, despite its more knowledge-intensive criteria.
-We hypothesize that this is due to differences in the labeling process—the n2c2 challenge involves more subjective judgment, while our setup explicitly defines insufficient information as ”NOT MET”.
-4.5.2. Error Analysis
-Despite MAKAR achieving 100% accuracy in the second experiment, we manually checked the reasoning results of all 300 patient-criterion pair and identified two reasoning errors that did not affect the final output.
-First, in Exclusion Criterion 13 (Patients with ongoing indications for cardioprotective medication—ACE inhibitors, ARBs, and/or beta-blockers), MAKAR reasoned: ”The patient is on metoprolol, a beta-blocker, suggesting some ongoing indication for its use.
-However, without explicit information on the specific condition necessitating this medication (e.g., arrhythmia, prophylactic use post-cardiac event other than MI, or another reason), the justification remains uncertain.” This reasoning was overly conservative, but the Supervision Matching Agent correctly classified it as MET, aligning with expert judgment.
-Second, in Inclusion Criterion 5 (Histologically confirmed breast cancer with an invasive component $\geq$ 20 mm and/or morphologically confirmed spread to regional lymph nodes(stage cT2-cT4 with any cN, or cN1-cN3 with any cT).), MAKAR reasoned: ”The documentation mentions a suspicious breast nodule at the right breast (1.4 × 1.5 × 0.5 cm), which measures more than 20 mm in one dimension.” This was an obvious error, as the maximum dimension stated is only 1.5 cm.
-The mistake may come from over-reasoning, possibly computing the largest possible diagonal as $\sqrt{1.4^{2}+1.5^{2}+0.5^{2}}=2.1cm$.
-However, the final decision remained correct since the criterion lymph node involvement still holds.
-4.6. Ablation
-In this section, we validate the contribution of each module in our framework.
-First, we remove the knowledge augmentation module to assess the reasoning module (MAKA-reason) using the original criteria, which essentially mirrors the CoT approach in Table 3.
-Next, we remove the reasoning module and evaluate the knowledge augmentation module (MAKA-aug) independently by performing zero-shot matching with the augmented criteria.
-The results, presented in Table 4, indicate that removing the reasoning module has a greater negative impact on inclusion criteria matching (F1-score: 88.37% vs.
-92.94%), while removing the augmentation module has a larger effect on exclusion criteria (F1-score: 93.65% vs.
-94.41%).
-Although we did not involve any reasoning-LLM-based agents, our approach still shows impressive reasoning capabilities.
-This can be attributed to the well-structured augmented criteria and the CoT strategy employed by the reasoning agent.
-For instance, when evaluating whether a patient meets the Inclusion Criterion 14 (Postmenopausal hormone receptor-positive patients), MAKA reasoned: ”…the record does not explicitly state whether the breast cancer was hormone receptor positive.
-However, the use of tamoxifen, which is specifically prescribed for estrogen receptor-positive breast cancer by blocking estrogen receptors, strongly implies that her breast cancer was hormone receptor positive.” Both MAKA and MAKA-reason successfully made the correct match, while MAKA-aug failed to capture this implicit relationship.
+To evaluate this, we manually reviewed the reasoning outputs for all 300 patient–criterion pairs.
+Although MAKAR produced entirely correct final classifications on the ClinicalTrial dataset, we identified two minor flaws in the intermediate reasoning steps.
+And the results reported in Table 1 reflect the quality of both the reasoning process and the final matching outcomes.
+First, in Exclusion Criterion 13 (Patients with ongoing indications for cardioprotective medication—ACE inhibitors, ARBs, and/or beta-blockers), MAKAR reasoned: "The patient is on metoprolol, a beta-blocker, suggesting some ongoing indication for its use.
+However, without explicit information on the specific condition necessitating this medication (e.g., arrhythmia, prophylactic use post-cardiac event other than MI, or another reason), the justification remains uncertain.".
+Here, the Reasoning Agent is overly conservative, but the Supervision Matching Agent correctly classified this case as MET.
+This case shows the design is effective.
+Second,in Inclusion Criterion 5 (Histologically confirmed breast cancer with an invasive component $\geq$ 20 mm and/or morphologically confirmed spread to regional lymph nodes(stage cT2-cT4 with any cN, or cN1-cN3 with any cT).), MAKAR wrongly inferred that a 1.5 cm lesion is greater than 20 mm.
+However, because lymph node involvement was documented, the final decision remained correct.
+Ablations Zero Shot MAKAR-Reason MAKAR-Aug MAKAR
+F1-Score $0.857$ $0.928$ $0.933$ $0.982$
+Reasoning Models
+OpenAI o1 $0.898$
+OpenAI o3 mini $0.825$
+Table 2: Results of Ablation Analysis. We report only zero-shot results for reasoning models here, as they are not suited for CoT strategies Guo et al. (2025).
+6 Ablation
+6.1 Augmentation and Reasoning
+In this section, we validate the double-module design of our workflow on ClinicalTrial dataset.
+First, we remove the knowledge augmentation module to assess the reasoning module alone (MAKAR-Reason, equivalent to a CoT) using the original criteria described in Appendix B.1.
+Next, we remove the reasoning module and evaluate the knowledge augmentation module (MAKAR-Aug) using zero-shot matching with prompts detailed in Appendix C.1).
+The results, presented in Table 2, suggest that both modules contributes to overall performance.
+Specifically, removing the augmentation module (MAKAR-Reason alone) led to a 6% drop in F1 score, while removing the reasoning module (MAKAR-Aug alone) led to a 5% drop.
+To further demonstrate the effectiveness of the reasoning module, we compared its performance against other reasoning models.
+Interestingly, MAKAR-Reason outperformed both OpenAI o1 and o3 mini, which can be attributed to the chain-of-thought strategy employed by the reasoning agent.
+For example, when evaluating whether a patient met Inclusion Criterion 14 (Postmenopausal hormone receptor-positive patients), MAKAR reasoned: "…the record does not explicitly state whether the breast cancer was hormone receptor positive.
+However, the use of tamoxifen, which is specifically prescribed for estrogen receptor-positive breast cancer by blocking estrogen receptors, strongly implies that her breast cancer was hormone receptor positive." Both MAKAR and MAKAR-Reason correctly identified this case, while MAKAR-Aug failed to capture the implicit relationship.
 Additionally, we also observed cases where knowledge augmentation improved performance.
-For example, in Exclusion Criterion 7, it mentions: ”…all histological lesions were HER2 1+ or 2+, not detected by FISH amplification, and HR positive according to ASCO guidelines”.
+For example, in Exclusion Criterion 7, it mentions: "…all histological lesions were HER2 1+ or 2+, not detected by FISH amplification, and HR positive according to ASCO guidelines".
 MAKA-aug expanded the definitions from American Society of Clinical Oncology(ASCO) guidelines, made correct matches for all 10 patients.
 In contrast, MAKA-reason correctly identified only 8 out of 10.
-Inclusion Criteria
-Metric MAKAR-reason MAKAR-aug MAKAR
-Accuracy 0.9667 0.9333 1.000
-Precision 0.9385 0.8387 1.000
-Recall 0.9208 0.9612 1.000
-F1-Score 0.9294 0.8837 1.000
-Exclusion Criteria
-Metric MAKAR-reason MAKAR-aug MAKARCorr
-Accuracy 0.9467 0.9533 1.000
-Precision 0.9265 0.9360 1.000
-Recall 0.9487 0.9534 1.000
-F1-Score 0.9365 0.9441 1.000
-Table 4. Results of Ablation Analysis
-5. Discussion
-One interesting observation from our experiments is that the Router Agent consistently determined that all criteria could be augmented by the Self-Augmentation Agent, without requiring external retrieval or online search augmentation.
+Underlying Models Zero Shot GPT-Aug MAKAR
+Qwen 2.5 Instruct 7B $0.853$ $0.853$ $0.900$
+Qwen 2.5 Instruct 14B $0.869$ $0.900$ $0.915$
+LLaMA 3.1 Instruct 7B $0.815$ $0.815$ $0.853$
+Gemma 3 Instruct 4B $0.788$ $0.788$ $0.828$
+Gemma 3 Instruct 12B $0.860$ $0.869$ $0.884$
+Mistral Instruct v0.3 7B $0.770$ $0.770$ $0.870$
+GPT4o 0.900 0.943 0.987
+Table 3: Results of Model Ablation in the Reasoning Module. GPT-Aug based on GPT-4o is equivalent to MAKAR-Aug.
+6.2 Local Deployment
+One key challenge for applying MAKAR in real-world settings is that agents in the reasoning module process patient EHR data.
+And such data is extremely sensitive and protected under The Health Insurance Portability and Accountability Act (HIPAA).
+It requires de-identifying large volumes of patient information before uploading data to a closed-source external model like GPT-4o.
+This process can be time-consuming and still have residual privacy risks.
+To address this limitation, we evaluated whether MAKAR can maintain high performance when replacing GPT-4o with open-source language models that can be deployed locally.
+Our local deployment experiment consisted of three configurations: (1)Zero-Shot Matching, where we applied open-source models directly in zero-shot mode using the criteria described in Appendix B.4 and the prompts from Appendix C.1; (2) GPT-Aug, where we used GPT-4 to augment the criteria (since criteria text itself is not sensitive) and then applied the same zero-shot prompt with a open-source model for matching; and (3) MAKAR, where we replaced all agents in the reasoning module with open-source models to avoid uploading data to externally hosted servers.
+The experiments were conducted on the ClinicalTrial dataset, and accuracies are reported in Table 3.
+We found that replacing GPT-4o in the reasoning module with locally deployed open-source models reduced performance by an average of 11%.
+However, some configurations still have competitive performance; for example, MAKAR powered by the Qwen family achieved accuracy greater than 0.900.
+These results further confirm the effectiveness of the MAKAR design, which improved performance by approximately 5% across open-source models.
+Another interesting observation is that smaller models appeared less sensitive to augmented criteria: models under 7B parameters showed no performance gains from augmentation.
+In contrast, both smaller and larger models consistently benefited from the reasoning module, with accuracy improving after reasoning was applied.
+However, this observation should be interpreted with caution, as it is based on a limited number of model configurations evaluated in this study.
+7 Related Work
+This work falls within the literature on patient matching using free-text patient records.
+Early approaches such as RBC Oleynik et al.
+(2019) relied on rule-based classification.
+Transformer encoder-based methods were later introduced, including DeepEnroll Zhang et al.
+(2020) and COMPOSE Gao et al.
+(2020), which improved the ability to encode complex eligibility criteria and patient notes.
+More recently, generative large language models (GLLMs) have been applied to this task and have shown stronger performance, as in TrialGPT Jin et al.
+(2024), CoT matching Beattie et al.
+(2024), and other zero-shot methods Wornow et al.
+(2024).
+In parallel, Loaiza-Bonilla et al.
+(2025) explored the use of multi-agent systems to analyze oncology knowledge graphs for patient enrollment.
+8 Conclusion
+In this study, we introduced MAKAR, a scalable multi-agent workflow designed to enhance patient-trial matching by combining criterion augmentation with structured reasoning.
+Our experimental results show that MAKAR achieves an average improvement of 7% and up to 10% on certain criteria.
+Through detailed ablation studies, we further demonstrate that both modules of the MAKAR design contribute independently to these gains.
+In addition, our local deployment experiments confirmed that MAKAR remains effective when using open-source language models.
+Even with smaller models, such as Qwen and Gemma, the MAKAR workflow achieved competitive performance.
+These findings suggest that MAKAR can play a significant role in enabling more accurate and privacy-preserving patient selection.
+Limitations
+One observation from our experiments is that the Router Agent consistently determined that all criteria could be augmented by the Self-Augmentation Agent, without requiring external retrieval or online search augmentation.
 This was true even for breast cancer-related criteria, which are typically complex and challenging for individuals without professional training.
@@
 Both the n2c2 dataset and the real-world dataset are relatively small and contain criteria that are not particularly challenging for LLM-based processing, allowing all benchmark methods to achieve strong performance.
-While MAKAR is designed as a general framework capable of handling both simple and complex criteria, these dataset constraints limit our ability to explore the knowledge boundaries of LLMs and fully assess the framework’s generalizability across a wider range of clinical trial requirements.
-Moreover, this observation raises a discussion about the specific improvements our framework brings to baseline LLMs.
-While the framework is designed to bridge the knowledge gap between trial criteria and LLMs’ existing knowledge, our experiments indicate that LLMs can already expand and interpret the criteria in our datasets accurately.
-Instead, the primary benefit of knowledge augmentation lies in explicitly organizing relevant concepts and definitions in a well-structured format.
-By clearly outlining these key elements, the augmentation provides a type of breakdown of the subsequent reasoning steps, leading to improved overall performance.
+While MAKAR is designed as a general framework capable of handling both simple and complex criteria, these dataset constraints limit our ability to explore the knowledge boundaries of LLMs and fully assess our method’s generalizability across a wider range of clinical trial requirements.
 However, despite the dataset limitations, the effectiveness of the augmentation module remains significant.
@@
 By systematically incorporating external knowledge, the augmentation module enhances the contextual relevance of trial matching, making it particularly valuable in real-world applications where timely and comprehensive knowledge retrieval is essential.
-6. Conclusion
-In this study, we introduced MAKAR, a scalable multi-agent framework designed to enhance patient-trial matching by combining knowledge augmentation with structured reasoning.
-By integrating an automated augmentation module with a reasoning module, MAKAR refines complex trial criteria and systematically evaluates patient eligibility.
-Our experimental results demonstrate that MAKAR outperforms existing benchmarks, including the state-of-the-art closed-source reasoning LLMs, and achieve 100% accuracy in a real-world offline test.
-And it also achieves up to an 8% improvement on complex criteria in a public dataset.
-Beyond performance gains, our findings emphasize the importance of explicitly structuring and expanding trial criteria to bridge the knowledge gap in AI-driven clinical reasoning.
-It is worth noting that the design of MAKAR supports the future development of a human-AI collaboration system.
-For the trial-centric patient matching task, the workload of criterion-supervision performed by the critic agent is one-time per trial and relatively small and can be efficiently handled by human experts with minimal time investment.
-Exploring how humans can further enhance and facilitate this system presents an interesting direction for future research.
-To address the limitations of this study and validate MAKAR’s generalizability, future work could also focus on evaluating the framework using more challenging datasets, such as those involving newly launched oncology trial that required up-to-date medical knowledge.
-Testing MAKAR in thess contexts would not only provide a more rigorous assessment of its capabilities but also demonstrate its potential to address critical real-world challenges in clinical trial recruitment.
-7. Acknowledgments
-We gratefully acknowledge support from the OpenAI Research Access Program, which provided access to advanced language models that were instrumental in developing and evaluating the framework presented in this work.
-Appendix A Criteria
-A.1. N2C2 Criteria Statistics
+Appendix A Statistics
 Criterion Met Count Not Met Count
@@
 MI-6MOS 26 262
-Table 5. Counts of Eligible and Ineligible Patients for Each Criterion
-A.2. N2C2 Original Criteria
-Criteria Definition
-DRUG-ABUSE Drug abuse, current or past.
-ALCOHOL-ABUSE Current alcohol use over weekly recommended limits.
-ENGLISH Patient must speak English.
-MAKES-DECISIONS Patient must make their own medical decisions.
-ABDOMINAL History of intra-abdominal surgery, small or large intestine resection, or small bowel obstruction.
-MAJOR-DIABETES Major diabetes-related complications, such as: amputation, kidney damage, skin conditions, retinopathy, nephropathy, neuropathy.
-ADVANCED-CAD Advanced cardiovascular disease, defined as two or more of the following: medications for CAD, history of myocardial infarction, angina, ischemia.
-MI-6MOS Myocardial infarction in the past 6 months.
-KETO-1YR Diagnosis of ketoacidosis in the past year.
-DIETSUPP-2MOS Taken a dietary supplement (excluding Vitamin D) in the past 2 months.
-ASP-FOR-MI Use of aspirin to prevent myocardial infarction.
-HBA1C Any HbA1c value between 6.5 and 9.5%.
-CREATININE Serum creatinine ¿ upper limit of normal.
-Table 6. Definitions of the original criteria in the n2c2 dataset.
-A.3. Redefined N2C2 Criteria
-Table 7. Redefined Criteria proposed by (Wornow et al., 2024)
-Criteria Definition
-ABDOMINAL History of intra-abdominal surgery. This could include any form of intra-abdominal surgery, including but not limited to small/large intestine resection or small bowel obstruction.
-ADVANCED-CAD Advanced cardiovascular disease (CAD).
-Defined as having 2 or more of the following: (a) Taking 2 or more medications to treat CAD (b) History of myocardial infarction (MI) (c) Currently experiencing angina (d) Ischemia, past or present.
+Table 4: Counts of Eligible and Ineligible Patients in N2C2
+Appendix B Criteria
+B.1 N2C2 Original Criteria
+DRUG-ABUSE: Drug abuse, current or past.
+ALCOHOL-ABUSE: Current alcohol use over weekly recommended limits.
+ENGLISH: Patient must speak English.
+MAKES-DECISIONS: Patient must make their own medical decisions.
+ABDOMINAL: History of intra-abdominal surgery, small or large intestine resection, or small bowel obstruction.
+MAJOR-DIABETES: Major diabetes-related complications, such as: amputation, kidney damage, skin conditions, retinopathy, nephropathy, neuropathy.
+ADVANCED-CAD: Advanced cardiovascular disease, defined as two or more of the following: medications for CAD, history of myocardial infarction, angina, ischemia.
+MI-6MOS: Myocardial infarction in the past 6 months.
+KETO-1YR: Diagnosis of ketoacidosis in the past year.
+DIETSUPP-2MOS: Taken a dietary supplement (excluding Vitamin D) in the past 2 months.
+ASP-FOR-MI: Use of aspirin to prevent myocardial infarction.
+HBA1C: Any HbA1c value between 6.5 and 9.5%.
+CREATININE: Serum creatinine > upper limit of normal.
+B.2 N2C2 Criteria Refined by Wornow et al. (2024)
+ABDOMINAL: History of intra-abdominal surgery. This could include any form of intra-abdominal surgery, including but not limited to small/large intestine resection or small bowel obstruction.
+ADVANCED-CAD: Advanced cardiovascular disease (CAD).
+For the purposes of this annotation, we define “advanced” as having 2 or more of the following: (a) Taking 2 or more medications to treat CAD (b) History of myocardial infarction (MI) (c) Currently experiencing angina (d) Ischemia, past or present.
 The patient must have at least 2 of these categories (a,b,c,d) to meet this criterion, otherwise the patient does not meet this criterion.
-For ADVANCED-CAD, be strict in your evaluation of the patient – if they just have cardiovascular disease, then they do not meet this criterion.
-ALCOHOL-ABUSE Current alcohol use over weekly recommended limits.
-ASP-FOR-MI Use of aspirin for preventing myocardial infarction (MI).
-CREATININE Serum creatinine level above the upper normal limit.
-DIETSUPP-2MOS Consumption of a dietary supplement (excluding vitamin D) in the past 2 months.
+For ADVANCED-CAD, be strict in your evaluation of the patient—if they just have cardiovascular disease, then they do not meet this criterion.
+ALCOHOL-ABUSE: Current alcohol use over weekly recommended limits.
+ASP-FOR-MI: Use of aspirin for preventing myocardial infarction (MI).
+CREATININE: Serum creatinine level above the upper normal limit.
+DIETSUPP-2MOS: Consumption of a dietary supplement (excluding vitamin D) in the past 2 months.
 To assess this criterion, go through the list of medications and supplements taken from the note.
@@
 Be lenient and broad in what is considered a dietary supplement.
-For example, a ’multivitamin’ and ’calcium carbonate’ should always be considered a dietary supplement if they are included in this list.
-DRUG-ABUSE Current or past history of drug abuse.
-ENGLISH Patient speaks English.
+For example, a “multivitamin” and “calcium carbonate” should always be considered a dietary supplement if they are included in this list.
+DRUG-ABUSE: Current or past history of drug abuse.
+ENGLISH: Patient speaks English.
 Assume that the patient speaks English, unless otherwise explicitly noted.
-If the patient’s language is not mentioned in the note, then assume they speak English and thus meet this criteria.
-HBA1C Any hemoglobin A1c (HbA1c) value between 6.5% and 9.5%.
-KETO-1YR Diagnosis of ketoacidosis within the past year.
-MAJOR-DIABETES Major diabetes-related complication.
-Examples of “major complication” (as opposed to “minor complication”) include, but are not limited to, any of the following that are a result of (or strongly correlated with) uncontrolled diabetes: • Amputation • Kidney damage • Skin conditions • Retinopathy • Nephropathy • Neuropathy.
+If the patient’s language is not mentioned in the note, then assume they speak English and thus meet this criterion.
+HBA1C: Any hemoglobin A1c (HbA1c) value between 6.5% and 9.5%.
+KETO-1YR: Diagnosis of ketoacidosis within the past year.
+MAJOR-DIABETES: Major diabetes-related complication.
+Examples of “major complication” (as opposed to “minor complication”) include, but are not limited to, any of the following that are a result of (or strongly correlated with) uncontrolled diabetes: Amputation, kidney damage, skin conditions, retinopathy, nephropathy, neuropathy.
 Additionally, if multiple conditions together imply a severe case of diabetes, then count that as a major complication.
-MAKES-DECISIONS Patient must make their own medical decisions.
+MAKES-DECISIONS: Patient must make their own medical decisions.
 Assume that the patient makes their own medical decisions, unless otherwise explicitly noted.
-There is no information provided about the patient’s ability to make their own medical decisions, then assume they do make their own decisions and therefore meet this criteria.
-MI-6MOS Myocardial infarction (MI) within the past 6 months.
-A.4. MAKAR-Augemented N2C2 Criteria
-Table 8. Redefined Criteria proposed by (Wornow et al., 2024)
-Criteria Definition
-ABDOMINAL Criteria: History of intra-abdominal surgery, including small/large intestine resection or small bowel obstruction.
+If there is no information provided about the patient’s ability to make their own medical decisions, then assume they do and therefore meet this criterion.
+MI-6MOS: Myocardial infarction (MI) within the past 6 months.
+B.3 N2C2 Criteria Refined by GPT-4o
+ABDOMINAL
+Criteria: History of intra-abdominal surgery, including small/large intestine resection or small bowel obstruction.
 Explanation: Intra-abdominal surgery involves operations on organs within the abdominal cavity (e.g., stomach, intestines).
 Small bowel obstruction refers to blockages in the small intestine that may require surgery.
-ADVANCED-CAD Criteria: Advanced cardiovascular disease (CAD) defined by having two or more of the following: (a) Taking 2 or more medications for CAD, (b) History of myocardial infarction (MI), (c) Current angina, or (d) Ischemia (past or present).
+ADVANCED-CAD
+Criteria: Advanced cardiovascular disease (CAD) defined by having two or more of the following: (a) Taking 2 or more medications for CAD, (b) History of myocardial infarction (MI), (c) Current angina, or (d) Ischemia (past or present).
 Explanation: CAD affects heart function due to blocked arteries.
@@
 A diagnosis of CAD alone does not qualify as advanced.
-ALCOHOL-ABUSE Criteria: Current alcohol use exceeding weekly recommended limits.
-Explanation: Weekly limits: up to 14 standard drinks. A standard drink is  14g of pure alcohol (12oz beer, 5oz wine, or 1.5oz spirits).
-ASP-FOR-MI Criteria: Use of aspirin to prevent myocardial infarction (MI).
+ALCOHOL-ABUSE
+Criteria: Current alcohol use exceeding weekly recommended limits.
+Explanation: Weekly limits: up to 14 standard drinks. A standard drink is approximately 14g of pure alcohol (12oz beer, 5oz wine, or 1.5oz spirits).
+ASP-FOR-MI
+Criteria: Use of aspirin to prevent myocardial infarction (MI).
 Explanation: Aspirin reduces MI risk by preventing clots but requires medical advice to balance benefits (risk reduction) and potential side effects (e.g., bleeding).
-CREATININE Criteria: Serum creatinine level above the upper normal limit.
-Explanation: Creatinine reflects kidney function. Normal levels vary by age and gender (e.g.,  0.6-1.2 mg/dL for adult males). Elevated levels suggest potential kidney impairment.
-DIETSUPP-2MOS Criteria: Consumption of a dietary supplement (excluding Vitamin D) in the past 2 months.
+CREATININE
+Criteria: Serum creatinine level above the upper normal limit.
+Explanation: Creatinine reflects kidney function. Normal levels vary by age and gender (e.g., approximately 0.6–1.2 mg/dL for adult males). Elevated levels suggest potential kidney impairment.
+DIETSUPP-2MOS
+Criteria: Consumption of a dietary supplement (excluding Vitamin D) in the past 2 months.
 Explanation: Supplements include multivitamins, calcium carbonate, fish oil, or herbal supplements like ginseng. Excludes Vitamin D explicitly.
-DRUG-ABUSE Criteria: Current or past drug abuse history.
+DRUG-ABUSE
+Criteria: Current or past drug abuse history.
 Explanation: Drug abuse includes misuse of psychoactive substances, with potential consequences like health issues or legal problems.
-ENGLISH Criteria: Patient must speak English.
+ENGLISH
+Criteria: Patient must speak English.
 Explanation: English proficiency ensures effective participation in English-based trials. Assume English proficiency unless noted otherwise.
-HBA1C Criteria: Hemoglobin A1c (HbA1c) between 6.5% and 9.5%.
+HBA1C
+Criteria: Hemoglobin A1c (HbA1c) between 6.5% and 9.5%.
 Explanation: HbA1c measures average blood sugar over 2–3 months. The range identifies diabetes patients with moderate control, excluding poorly controlled cases.
-KETO-1YR Criteria: Diagnosis of ketoacidosis in the past year.
+KETO-1YR
+Criteria: Diagnosis of ketoacidosis in the past year.
 Explanation: Ketoacidosis is a serious condition with high blood ketones, often related to uncontrolled diabetes. Symptoms include nausea, confusion, and rapid breathing.
-MAJOR-DIABETES Criteria: Major diabetes-related complications like amputation, kidney damage, or neuropathy.
+MAJOR-DIABETES
+Criteria: Major diabetes-related complications like amputation, kidney damage, or neuropathy.
 Explanation: Includes significant health effects (e.g., nephropathy, retinopathy) stemming from uncontrolled diabetes. Multiple minor complications combined may also qualify.
-MAKES-DECISIONS Criteria: Patient must make their own medical decisions.
+MAKES-DECISIONS
+Criteria: Patient must make their own medical decisions.
 Explanation: This assesses cognitive ability to understand medical options, communicate choices, and appreciate consequences.
-MI-6MOS Criteria: Myocardial infarction (MI) within the past 6 months.
+MI-6MOS
+Criteria: Myocardial infarction (MI) within the past 6 months.
 Explanation: MI refers to heart muscle damage from blocked blood flow. Assess via medical history and records confirming the event within 6 months.
-A.5. Sampled ClinicalTrial.gov Criteria
-Inclusion Criteria Exclusion Criteria
-1. Locally advanced or metastatic breast cancer confirmed by histopathology. 1. Patients known or suspected intolerance or hypersensitivity to main ingredient or any of the excipients of SNB-101.
-2. Received 2 previous lines of anti-tumor treatment and developed resistance to standard treatment. 2. Patients with bilateral invasive breast cancers.
+B.4 Criteria from ClinicalTrials.gov
+Inclusion Criteria
+1. Locally advanced or metastatic breast cancer confirmed by histopathology.
+2. Received 2 previous lines of anti-tumor treatment and developed resistance to standard treatment.
 3.
 Patients with locally advanced or metastatic advanced solid tumors confirmed by histology or cytology (including but not limited to triple-negative breast cancer, gastric cancer, colorectal cancer); Patients with 1 line of standard treatment failure (disease progression after treatment or intolerability of toxic side effects of treatment), or no standard treatment, or unable to receive standard treatment.
-3.
-Age $<$18 years old or $>$70 years old.
 4.
 At least 1 measurable lesion per Response Evaluation Criteria in Solid Tumors version 1.1 and previously treated lesions with radiotherapy or focal therapy and no progression cannot be included as target lesion for assessment.
-4.
-Patients with standard metallic contraindications to CMR or estimated glomerular filtration rate $<$ 30 mL/min/1.73 m².
 5.
 Histologically confirmed breast cancer with an invasive component measuring 20 mm and/or with morphologically confirmed spread to regional lymph nodes (stage cT2-cT4 with any cN, or cN1-cN3 with any cT).
-5.
-History of hypersensitivity or contraindication to TTF.
-6.
-Known HER2-positive breast cancer defined as an IHC status of 3+.
-If IHC is 2+, a positive in situ hybridization (FISH, CISH, or SISH) test is required by local laboratory testing.
-6.
-Implanted pacemaker, defibrillator, or other electrical medical devices.
-7.
-Patients with breast cancer planned to receive Anthracycline and Cyclophosphamide chemotherapy.
-7.
-Bilateral breast cancer (including multifocal breast cancer.
-All histological lesions were HER2 1+ or 2+, not detected by FISH amplification and HR positive according to ASCO guidelines).
-8.
-Patients treated for hormone-dependent localized breast cancer requiring adjuvant hormonal therapy (HT).
-8.
-Received any form of anti-tumor therapy (chemotherapy, radiotherapy, molecular targeted therapy, endocrine therapy, etc.).
-9. Patients treated with tamoxifen for a maximum of 1 to 3 years. 9. Breast cancer without histopathological diagnosis.
-10.
-Histologically confirmed ER+/HER2- early-stage resected invasive breast cancer at high or intermediate risk of recurrence, , based on clinical-pathological risk features, as defined in the protocol.
-10.
-Known personal history of ductal carcinoma in situ (DCIS) or invasive breast cancer.
-11.
-Completed adequate (definitive) locoregional therapy (surgery with or without radiotherapy) for the primary breast tumour(s), with or without (neo)adjuvant chemotherapy.
-11.
-Prior systemic treatment for any malignancy.
-12. Completed at least 2 years but no more than 5 years (+3 months) of adjuvant ET (+/- CDK4/6 inhibitor). 12. Significant medical comorbidities as per investigator evaluation.
-13. Patients receiving letrozole for more than two months. 13. Patients with ongoing indications for the cardioprotective medication - ACE inhibitors, ARBs and/or beta-blockers
-14. Postmenopausal hormone receptor-positive patients. 14. Intermediate or high-grade ductal carcinoma in situ.
-15. Patients with N0 disease or those who have undergone targeted axillary detection with a good response. 15. Invasive carcinoma.
-Table 9. Sampled Criteria from Breast Cancer Clinical Trials
+6. Known HER2-positive breast cancer defined as an IHC status of 3+. If IHC is 2+, a positive in situ hybridization (FISH, CISH, or SISH) test is required by local laboratory testing.
+7. Patients with breast cancer planned to receive Anthracycline and Cyclophosphamide chemotherapy.
+8. Patients treated for hormone-dependent localized breast cancer requiring adjuvant hormonal therapy (HT).
+9. Patients treated with tamoxifen for a maximum of 1 to 3 years.
+10. Histologically confirmed ER+/HER2- early-stage resected invasive breast cancer at high or intermediate risk of recurrence, based on clinical-pathological risk features, as defined in the protocol.
+11. Completed adequate (definitive) locoregional therapy (surgery with or without radiotherapy) for the primary breast tumour(s), with or without (neo)adjuvant chemotherapy.
+12. Completed at least 2 years but no more than 5 years (+3 months) of adjuvant ET (+/- CDK4/6 inhibitor).
+13. Patients receiving letrozole for more than two months.
+14. Postmenopausal hormone receptor-positive patients.
+15. Patients with N0 disease or those who have undergone targeted axillary detection with a good response.
+Exclusion Criteria
+1. Patients known or suspected intolerance or hypersensitivity to main ingredient or any of the excipients of SNB-101.
+2. Patients with bilateral invasive breast cancers.
+3. Age $<$18 years old or $>$70 years old.
+4. Patients with standard metallic contraindications to CMR or estimated glomerular filtration rate $<$30 mL/min/1.73 m².
+5. History of hypersensitivity or contraindication to TTF.
+6. Implanted pacemaker, defibrillator, or other electrical medical devices.
+7. Bilateral breast cancer (including multifocal breast cancer. All histological lesions were HER2 1+ or 2+, not detected by FISH amplification and HR positive according to ASCO guidelines).
+8. Received any form of anti-tumor therapy (chemotherapy, radiotherapy, molecular targeted therapy, endocrine therapy, etc.).
+9. Breast cancer without histopathological diagnosis.
+10. Known personal history of ductal carcinoma in situ (DCIS) or invasive breast cancer.
+11. Prior systemic treatment for any malignancy.
+12. Significant medical comorbidities as per investigator evaluation.
+13. Patients with ongoing indications for the cardioprotective medication - ACE inhibitors, ARBs and/or beta-blockers.
+14. Intermediate or high-grade ductal carcinoma in situ.
+15. Invasive carcinoma.
+B.5 Example Criteria from ClinicalTrials.gov refined by GPT-4o
+Criterion: Locally advanced or metastatic breast cancer confirmed by histopathology.
+Additional Explanation:
+- Locally advanced breast cancer refers to cancer that has spread beyond the breast to nearby areas but not to distant body parts. It often involves large tumors or extensive lymph node involvement.
+- Metastatic breast cancer means that the cancer has spread from the breast to other parts of the body, such as the bones, liver, lungs, or brain.
+- Histopathology is the examination of tissues under a microscope to study the manifestations of disease.
+This confirms the diagnosis of breast cancer by analyzing tissue samples obtained from a biopsy or surgery.
+Criterion: Received 2 previous lines of anti-tumor treatment and developed resistance to standard treatment.
+Additional Explanation:
+1. **Anti-tumor treatment**: This refers to any therapy aimed at treating cancer, which might include chemotherapy, targeted therapy, immunotherapy, or hormone therapy.
+2.
+**2 previous lines of treatment**: In cancer therapy, a "line" of treatment refers to a complete sequence of therapy given for cancer, which can consist of one or multiple medications or interventions.
+The patient must have undergone two different complete treatment sequences.
+3.
+**Developed resistance**: This means that the cancer no longer responds to the standard treatment.
+This can be defined by the tumor growing or not shrinking in size despite treatment, or the cancer markers take not being reduced.
+4.
+**Standard treatment**: This refers to the most widely accepted and utilized treatments for a particular type of cancer, as established by clinical research and guidelines.
+Examples might include first-line chemotherapy regimens or targeted therapies.
+Exact "standard treatment" could vary based on cancer type, but usually involves those strategies that have the highest success rates based on evidence.
+Appendix C Prompts
+C.1 The prompt for zero-shot matching
+You are an experienced physician tasked with determining whether a patient meets specific eligibility criteria for clinical trials.
+Below, I will provide multiple EHR records for a single patient along with one criterion.
+You must evaluate whether the patient satisfies the criterion.
+For each criterion, respond strictly with 1 if the criterion is met, or 0 if it is not met.
+If there is not enough information to make a determination, return 0.
+Respond with only 0 or 1; no reasoning or explanation is needed.
+Patient EHR Data:…
+The Ctiterion:…
+C.2 The prompt for refining criteria
+You are an experienced physician tasked with reviewing the trial criterion to make it easy to follow for less-trained staff.
+Sometimes, the criterion may seem vague.
+If it is, please complement the criterion for clarity in plain English by: (1) defining any specialized medical terms, and (2) explaining vague phrases such as ‘normal levels’ by specifying normal ranges for different genders and age groups.
+You must keep the original criterion intact and add additional explanations only if you think they are necessary.
+Please ONLY return the revised criterion and use the following format:
+<Original Criterion>: XXX
+<Additional Explanation>: XXX
+Here is the original criterion:…
+C.3 Prompts for Agents
+C.3.1 Router Agent
+You are a routing agent responsible for selecting the most appropriate augmentation agent to enrich a trial criterion with relevant, domain-specific information.
+You will receive one criterion as input.
+Based on its content, determine which augmentation agent to use.
+**Retrieval Agent**: Use this agent if the criterion requires authoritative information from indexed medical databases like MedDRA and FDALabel.
+**Online Search Agent**: Use this agent if the criterion needs supplementary knowledge from external online sources.
+**Self-Augment Agent**: Use this agent if the criterion can be refined using the language model’s intrinsic medical knowledge.
+Strictly respond with only one of the following labels indicating your choice: Retrieval, Online Search, or Self-Augment. No reasoning or explanation is needed—only return the label.
+The Criterion:..
+C.3.2 Knowledge Augmentation Agent
+You are an experienced physician tasked with reviewing the trial criterion to make it easy to follow for less-trained staff.
+Sometimes, the criterion may seem vague.
+If it is, please complement the criterion for clarity in plain English by: (1) defining any specialized medical terms, and (2) explaining vague phrases such as ‘normal levels’ by specifying normal ranges for different genders and age groups.
+You must keep the original criterion intact and add additional explanations only if you think they are necessary.
+Please ONLY return the revised criterion and use the following format:
+<Original Criterion>: XXX
+<Additional Explanation>: XXX
+Here is the original criterion:…
+C.3.3 Critic Agent
+You are a senior physician reviewing criteria revised by another physician.
+Make sure that (1) the original criterion is preserved, and (2) any additional explanations are accurate.
+For each criterion, respond strictly with **1** if both checks pass, or **0** if they do not.
+No reasoning or explanation is needed.
+The Revised Ctiterion:…
+The Original Ctiterion:…
+C.3.4 Reasoning Agent
+You are a Principal Investigator (PI) in a clinical trial.
+You have multiple EHR records for one patient and a criterion with detailed explanation.
+Please perform a reasoning process to analyze why this patient meets or does not meet this trial criterion.
+If the information is not sufficient to make a determination, assume the patient does not meet the criterion.
+Patient EHR Data:…
+The Detailed Ctiterion:…
+C.3.5 Matching Agent
+You are a Principal Investigator (PI) in a clinical trial and are tasked with determining whether a patient meets specific eligibility criteria.
+Below, I will provide multiple EHR records for one patient, one criterion, and a reasoning process from another PI.
+You must evaluate whether the patient meets the criterion.
+For each criterion, respond strictly with **1** if met or **0** if not met.
+Respond with 0 or 1 only; no reasoning or explanation is needed.
+Patient EHR Data:…
+The Ctiterion:…
+The Reasoning Process:…
+Appendix D Pseudo-Code for MAKAR Workflow
+Algorithm 1 MAKAR: Multi-Agent Knowledge Augmentation and Reasoning
+1:Original criterion $C_{orig}$, Patient EHR data $P_{EHR}$
+2:Matching result $M\in\{\text{0},\text{1}\}$
+3:// Augmentation Module: MAKAR-Aug
+4:$C_{aug}\leftarrow\text{RouterAgent}(C_{orig})$
+5:if RouterAgent determines $\text{RouterAgent determines }C_{orig}\text{ is detailed enough}$ then
+6:     $C_{aug}\leftarrow C_{orig}$triangleright Forward directly to Reasoning Module
+7:else
+[…]
+[…]
+10:         $C_{aug}\leftarrow\text{RetrievalAgent}(C_{orig},\text{databases})$
+[…]
+12:         $C_{aug}\leftarrow\text{OnlineSearchAgent}(C_{orig})$
+[…]
+14:         $C_{aug}\leftarrow\text{SelfAugmentAgent}(C_{orig})$
+15:     end if
+16:     // Iterative refinement with Critic Agent
+17:     repeat
+18:         $\text{feedback}\leftarrow\text{CriticAgent}(C_{aug},C_{orig})$
+19:         if feedback.rejected then
+20:              $C_{aug}\leftarrow\text{Re-augment}(C_{orig},\text{feedback})$
+21:         end if
+22:     until CriticAgent accepts $\text{CriticAgent accepts }C_{aug}$
+23:end if
+24:// Reasoning Module: MAKAR-Reason
+[…]
+26:// Generate detailed evaluation
+27:for each concept $c$ in $C_{aug}$ do
+28:     $\text{evaluation}[c]\leftarrow\text{ReasoningAgent.evaluate}(c,P_{EHR})$
+29:end for
+30:// Final matching decision
+[…]
+32:// Supervision validation
+[…]
+34:if validation.approved then
+35:     return $M$
+36:else
+37:     return validation.corrected_result
+38:end if